### v1

# On generalized Turán problems for expansions

Abstract
Given a graph $F$, the $r$-expansion $F^{r}$ of $F$ is the $r$-uniform hypergraph obtained from $F$ by inserting r2r-2 new distinct vertices in each edge of $F$. Given $r$-uniform hypergraphs $\mathcal{H}$} and $\mathcal{F}$}, the generalized Turán number, denoted by $\mathrm{ex}_{r}(n,{\mathcal{H}},{\mathcal{F}})$, is the maximum number of copies of $\mathcal{H}$} in an $n$-vertex $r$-uniform hypergraph that does not contain $\mathcal{F}$} as a subhypergraph. In the case where $r=2$ (i.e., the graph case), the study of generalized Turán problems was initiated by Alon and Shikhelman [J. Combin. Theory Series B. 121 (2016) 146–172]. Motivated by their work, we systematically study generalized Turán problems for expansions and obtain several general and exact results. In particular, for the non-degenerate case, we determine the exact generalized Turán number for expansions of complete graphs, and establish the asymptotics of the generalized Turán number for expansions of the vertex-disjoint union of complete graphs. For the degenerate case, we establish the asymptotics of generalized Turán numbers for expansions of several classes of forests, including star forests, linear forests and star-path forests.
footnotetext: *Corresponding author.footnotetext: Email addresses: junpengzhou@shu.edu.cn (J. Zhou), zxm23@mails.tsinghua.edu.cn (X. Zhao) xiyingyuan@shu.edu.cn (X. Yuan).
Keywords: hypergraph, generalized Turán problem, expansion
AMS subject classifications: 05C35, 05C65
1 Introduction
An $r$-uniform hypergraph ($r$-graph for short) ${\mathcal{H}}=(V({\mathcal{H}}),E({\mathcal{H}}))$ consists of a vertex set $V({\mathcal{H}})$ and a hyperedge set $E({\mathcal{H}})$, where each hyperedge in $E({\mathcal{H}})$ is an $r$-subset of $V({\mathcal{H}})$. The size of $E({\mathcal{H}})$ is denoted by $e({\mathcal{H}})$. Let ${\mathcal{H}}_{1}{\cup}{\mathcal{H}}_{2}$ denote the vertex-disjoint union of $r$-graphs ${\mathcal{H}}_{1}$ and ${\mathcal{H}}_{2}$, and let $k{\mathcal{H}}$ denote the vertex-disjoint union of $k$ $r$-graphs $\mathcal{H}$}.
Let $\mathcal{F}$ be an $r$-graph. An $r$-graph $\mathcal{H}$} is $\mathcal{F}$-free if $\mathcal{H}$} does not contain $\mathcal{F}$ as a subhypergraph. The Turán number of $\mathcal{F}$, denoted by ${\rm{ex}}_{r}(n,\mathcal{F})$, is the maximum number of hyperedges in an $n$-vertex $\mathcal{F}$-free $r$-graph. When $r=2$, we use $\mathrm{ex}(n,\mathcal{F})$ instead of $\mathrm{ex}_{2}(n,\mathcal{F})$, which represents a fundamental and well-studied problem in extremal graph theory. A classical result in extremal graph theory is Turán theorem [48], which determines the exact Turán number for the $\ell$-vertex complete graph $K_{\ell}$. The Erdős-Stone-Simonovits theorem [16, 17] gives an asymptotics of the Turán number for any $k$-chromatic graph. When $\mathcal{F}$} is bipartite, the problem of determining $\mathrm{ex}(n,{\mathcal{F}})$ remains an active topic in extremal graph theory. For an extensive overview of the historical development, we refer the reader to the survey by Füredi and Simonovits [27]. In particular, Erdős and Sós [13] conjectured ${\rm ex}(n,{\mathcal{F}})\leq\frac{n(\ell-1)}{2}$ when $\mathcal{F}$} is a tree with $\ell$ edges. The conjecture is known to hold for paths by the Erdős-Gallai theorem [15] and for spiders [18]. Unlike the conjecture proposed for trees, the Turán number of a forest is highly dependent on its structural characteristics [36]. Let $M_{s+1}$ denote a matching of size $s+1$, i.e., the graph consisting of $s+1$ independent edges. In 1959, Erdős and Gallai [15] determined the Turán number of $M_{s+1}$. Let $P_{\ell}$ and $S_{\ell}$ denote a path and a star with $\ell$ edges, respectively. A star forest is a forest whose connected components are stars and a linear forest is a forest whose connected components are paths. Bushaw and Kettle [8] determined the Turán number of $kP_{\ell}$ for $k\geq 2$, $\ell\geq 3$ and sufficiently large $n$. Later, Lidický, Liu and Palmer [40] generalized the result of Bushaw and Kettle [8], and simultaneously determined the Turán number of star forests.
It is well known that hypergraph Turán problems are considerably more difficult than in the graph setting. Some natural questions are surprisingly difficult. For example, the Turán number and even the asymptotics for $K_{4}^{(3)}$, the 4-vertex complete $3$-graph, remains unknown. The open problem shows that even for hypergraphs with few vertices and hyperedges, the Turán problem may be complicated. A fundamental object of investigation in hypergraph Turán problems is expansions.
Given a graph $F$, the $r$-uniform expansion (or briefly $r$-expansion) $F^{r}$ of $F$ is the $r$-graph obtained from $F$ by inserting r2r-2 new distinct vertices in each edge of $F$, such that the (r2)e(F)(r-2)e(F) new vertices are distinct from each other and are not in $V(F)$. It was introduced by Mubayi in [43]. Turán problems on expansions have been extensively studied, yielding numerous related extremal results. For expansions of complete graphs, Erdős [12] conjectured that ${\rm{ex}}_{r}(n,K_{3}^{r})=\binom{n-1}{r-1}$ for $n\geq\frac{3}{2}r$. The conjecture was later proved by Mubayi and Verstraëte [44]. Let ${\mathcal{T}}_{r}(n,\ell)$ denote the complete balanced $\ell$-partite $r$-graph. For convenience, let $t_{r}(n,\ell):=e({\mathcal{T}}_{r}(n,\ell))$. Mubayi [43] conjectured that ${\mathcal{T}}_{r}(n,\ell)$ is the unique maximum $K_{\ell+1}^{r}$-free $r$-graph for sufficiently large $n$. The conjecture was later proved by Pikhurko [46].
Theorem 1.1 (Pikhurko [46]).
Let integers $\ell\geq r\geq 3$. If $n$ is sufficiently large, then $\mathrm{ex}_{r}(n,K_{\ell+1}^{r})=t_{r}(n,\ell)$ and ${\mathcal{T}}_{r}(n,\ell)$ is the unique extremal hypergraph.
Gerbner [28] determined the exact Turán number for expansions of the vertex-disjoint union of complete graphs. Let $S_{n,t}^{r}(n-t,\ell)$ denote the following $r$-graph. We take a set $U$ of $t$ vertices and a copy of ${\mathcal{T}}_{r}(n-t,\ell)$, then add each $r$-set containing at least one vertex from $U$ as a hyperedge.
Theorem 1.2 (Gerbner [28]).
Let $F$ consist of $k$ components with chromatic number $\ell+1$, each with a color-critical edge, and any number of components with chromatic number at most $\ell$. Then ${\rm{ex}}_{r}(n,F^{r})=e(S_{n,k-1}^{r}(n-k+1,\ell))$ for sufficiently large $n$.
For expansions of stars, Duke and Erdős [10] gave an upper bound for ${\rm{ex}}_{r}(n,S_{\ell}^{r})$ as follows.
Theorem 1.3 (Duke and Erdős [10]).
Fix integers $r\geq 3$ and $\ell\geq 2$. Then there exists a constant $c(r)$ such that for sufficiently large $n$,
${\rm{ex}}_{r}(n,S_{\ell}^{r})\leq c(r)\ell(\ell-1)n^{r-2}.$
For expansions of paths, Füredi, Jiang and Seiver [26] determined ${\rm{ex}}_{r}(n,P_{\ell}^{r})$ for $r,\ell\geq 4$ and sufficiently large $n$. The case when $r\geq 3$ and $\ell\geq 4$ was proved by Kostochka, Mubayi and Verstraëte [38] using methods that differ significantly from those in [26]. The case $\ell=3$ was settled by Frankl and Füredi [23]. The case when $\ell=2$ and $r\geq 4$ was settled by Frankl [19] and the case when $\ell=2$, $r=3$ and all $n$ was settled by Sós [47]. Furthermore, Kostochka, Mubayi and Verstraëte [38] also determined ${\rm{ex}}_{r}(n,C_{\ell}^{r})$ for all $r\geq 3$, $\ell\geq 4$ and sufficiently large $n$, where $C_{\ell}$ denotes a cycle with $\ell$ edges.
Now let us consider expansions of forests. In 1965, Erdős [11] proposed the following conjecture for matchings.
Conjecture 1.4 (Erdős Matching Conjecture [11]).
Let integers $r\geq 2$, $s\geq 1$ and $n\geq(s+1)r-1$. Then
$\mathrm{ex}_{r}(n,M_{s+1}^{r})\leq\max\left\{\binom{(s+1)r-1}{r},\binom{n}{r}-\binom{n-s}{r}\right\}$.
It is worth noting that the conjecture is known to hold for the cases $r=2$ [15] and $r=3$ [20, 42]. For the general case, Erdős [11] verified Conjecture 1.4 for large $n$. Later, the threshold was improved in several papers by Bollobás, Daykin, Erdős [5], Huang, Loh, Sudakov [35], Frankl, Łuczak, Mieczkowska [25]. For further developments and relevant results concerning this conjecture, one can refer to [21, 24, 22, 37]. Very recently, Kupavskii and Sokolov [39] completely resolved the conjecture for the case $n\leq 3(s+1)$.
Bushaw and Kettle [7] determined the exact Turán number for expansions of linear forests when $n$ is sufficiently large, as follows.
Theorem 1.5 (Bushaw and Kettle [7]).
Let $r\geq 3$, $k\geq 2$, $\ell_{1},\ldots,\ell_{k}\geq 1$, and $n$ be sufficiently large. Setting $t=\sum_{i=1}^{k}\lfloor\frac{\ell_{i}+1}{2}\rfloor-1$ and $d_{\ell}=0$ when at least one of the $\ell_{i}$ is odd and $d_{\ell}=\dbinom{n-t-2}{r-2}$ otherwise, the following holds:
$\mathrm{ex}_{r}(n,P_{\ell_{1}}^{r}\cup\cdots\cup P_{\ell_{k}}^{r})=\dbinom{n-1}{r-1}+\cdots+\dbinom{n-t}{r-1}+d_{\ell}.$
Khormali and Palmer [36] obtained the following result for expansions of star forests.
Theorem 1.6 (Khormali and Palmer [36]).
Fix integers $\ell,k\geq 1$ and $r\geq 2$. Then for sufficiently large $n$,
${\rm{ex}}_{r}(n,kS_{\ell}^{r})=\binom{n}{r}-\binom{n-k+1}{r}+{\rm{ex}}_{r}(n-k+1,S_{\ell}^{r}).$
Recently, Zhou and Yuan [50, 51] studied the Turán problem on expansions of star-path forests, where a star-path forest is a forest whose connected components are either paths or stars. For further results on Turán problems for expansions, one can refer to Subsection 5.2.1 of [33], as well as the survey [45] by Mubayi and Verstraëte.
Given two graphs $H$ and $F$, let $\mathrm{ex}(n,H,F)$ denote the maximum number of copies of $H$ in an $n$-vertex $F$-free graph. Alon and Shikhelman [1] initiated the systematic study of the function $\mathrm{ex}(n,H,F)$, which is often called the generalized Turán problem. In the case $H=K_{s}$, Zykov [53] and independently Erdős [14] determined the exact value of $\mathrm{ex}(n,K_{s},K_{\ell})$. Wang [49] determined the exact value of $\mathrm{ex}(n,K_{s},M_{k+1})$. In [9], Chakraborti and Chen determined the exact value of $\mathrm{ex}(n,K_{s},P_{\ell})$. Given $r$-graphs $\mathcal{H}$} and $\mathcal{G}$}, let $\mathcal{N}({\mathcal{H}},{\mathcal{G}})$ denote the number of copies of $\mathcal{H}$} in $\mathcal{G}$}.
Theorem 1.7 (Chakraborti and Chen [9]).
For any positive integers $n$ and $3\leq s\leq\ell$, we have $\mathrm{ex}(n,K_{s},P_{\ell})={\mathcal{N}}(K_{s},pK_{\ell}\cup K_{q})$, where $n=p\ell+q$ and $0\leq q\leq\ell-1$.
Later, Zhu and Chen [52] determined $\mathrm{ex}(n,K_{s},F)$ when $F$ is the vertex-disjoint union of paths of length at least $4$ and $n$ is sufficiently large. Khormali and Palmer [36] determined $\text{ex}(n,K_{s},kS_{\ell})$ if $3\leq s\leq\ell+k-2$, $\ell$ divides nk+1n-k+1 and $n$ is sufficiently large. For a very recent survey on generalized Turán problems, one can refer to the work of Gerbner and Palmer [30].
It is natural to investigate the generalized Turán problem on hypergraphs. However, determining generalized Turán numbers for hypergraphs is more challenging than determining their classical Turán number counterparts. Nevertheless, several sporadic results on generalized Turán problems for hypergraphs have still been established. Before stating these results, we introduce several notations. Given $r$-graphs $\mathcal{H}$} and $\mathcal{F}$}, let $\mathrm{ex}_{r}(n,{\mathcal{H}},{\mathcal{F}})$ denote the maximum number of copies of $\mathcal{H}$} in an $n$-vertex $\mathcal{F}$}-free $r$-graph. When $r=2$, we use $\mathrm{ex}(n,{\mathcal{H}},{\mathcal{F}})$ instead of $\mathrm{ex}_{2}(n,{\mathcal{H}},{\mathcal{F}})$. We denote by $K_{s}^{(r)}$ the complete $r$-graph on $s$ vertices if $s\geq r$, and define $K_{s}^{(r)}$ as an $s$-element set if $s<r$. For $s<r$, let $\mathcal{N}(K_{s}^{(r)},{\mathcal{H}})$ denote the collection of all $s$-subsets of $V({\mathcal{H}})$. Let $n,k,r,a$ be positive integers with $n\geq r\geq a\geq 1$. Define the $r$-graph
$\mathcal{F}_{n,k,a}^{(r)}=\left\{F\in\binom{[n]}{r}:\big|F\cap[ak+a-1]\big|\geq a\right\}.$
Liu and Wang [41] determined $\text{ex}_{r}(n,K_{s}^{(r)},M^{r}_{k+1})$ for all $s,r,k$ and sufficiently large $n$.
Theorem 1.8 (Liu and Wang [41]).
Let $n,k,r,s$ be integers.
(i) If $r\leq s\leq k+r-1$ and $n\geq 4(er)^{s-r+2}k$, then $\mathrm{ex}_{r}(n,K_{s}^{(r)},M^{r}_{k+1})\leq{\mathcal{N}}(K_{s}^{(r)},\mathcal{F}_{n,k,1}^{(r)})$;
(ii) If $k+r\leq s\leq(r-1)(k+1)$ and $n\geq 4r^{2}k(er/(a-1))^{s-r+a}$, then $\mathrm{ex}_{r}(n,K_{s}^{(r)},M^{r}_{k+1})\leq{\mathcal{N}}(K_{s}^{(r)},\mathcal{F}_{n,k,a}^{(r)})$, where $a=\left\lfloor\frac{s-r}{k}\right\rfloor+1$;
(iii) If $(r-1)k+r\leq s\leq rk+r-1$ and $n\geq rk+r-1$, then $\mathrm{ex}_{r}(n,K_{s}^{(r)},M^{r}_{k+1})\leq{\mathcal{N}}(K_{s}^{(r)},\mathcal{F}_{n,k,r}^{(r)})$.
Note that $M^{r}_{k+1}=(k+1)K_{r}^{(r)}$. Gerbner [29] considered a generalization of Theorem 1.8, namely, the vertex-disjoint union of hypergraph cliques. We introduce several definitions from [29]. For two $r$-graphs $\mathcal{H}$ and $\mathcal{H}^{\prime}$ with disjoint vertex sets, we denote by $\mathcal{H}+\mathcal{H}^{\prime}$ the $r$-graph containing $\mathcal{H}$ and $\mathcal{H}^{\prime}$, that contains as additional hyperedges all the $r$-sets intersecting both $\mathcal{H}$ and $\mathcal{H}^{\prime}$. Let $x=\lceil\frac{kp-s}{k-1}\rceil-1$ and let $\mathcal{T}$ be a $K_{x+1}^{(r)}$-free $r$-graph on nk(px)+1n-k(p-x)+1 vertices with the most copies of $K_{x}^{(r)}$.
Theorem 1.9 (Gerbner [29]).
Let $s\geq p\geq 2$ and $k\geq 1$ be arbitrary integers and let $x=\lceil\frac{kp-s}{k-1}\rceil-1$. Then we have $\mathrm{ex}_{r}(n,K_{s}^{(r)},kK_{p}^{(r)})=(1+o(1))\mathcal{N}(K_{s}^{(r)},K_{k(p-x)-1}^{(r)}+\mathcal{T})$. In particular, $\mathrm{ex}_{r}(n,K_{s}^{(r)},kK_{p}^{(r)})=\Theta(n^{x})$.
Very recently, Axenovich, Gerbner, Liu and Patkós [2] established a generalized hypergraph Turán result for expansions of paths and cycles. For integers $r\geq 2$ and $t\geq 1$, let $S_{n,t}^{r}$ be the $n$-vertex $r$-graph consisting of all hyperedges that intersect a fixed set of $t$ vertices.
Theorem 1.10 (Axenovich, Gerbner, Liu and Patkós [2]).
Let $r\geq 3$ and $\ell\geq 4$ be integers and $\ell^{\prime}=\lfloor\frac{\ell-1}{2}\rfloor$. Let $Q\in\{P_{\ell}^{r},C_{\ell}^{r}\}$.
(i) If $s\leq\ell^{\prime}+r-1$, then $\mathrm{ex}_{r}(n,K_{s}^{(r)},Q)=(1+o(1))\mathcal{N}(K_{s}^{(r)},S_{n,\ell^{\prime}}^{r})$.
(ii) If $s\geq\ell^{\prime}+r$, then $\mathrm{ex}_{r}(n,K_{s}^{(r)},Q)=o(n^{r-1})$.
To the best of our knowledge, besides the results mentioned above, the only other generalized Turán result on hypergraphs is that concerning complete multipartite $r$-graphs, due to Basu, Rödl and Zhao [3]. Specifically, they showed that for $r\geq 2$ and $a_{1}\leq\cdots\leq a_{r+1}$,
$\mathrm{ex}_{r}\big(n,K_{r+1}^{(r)},K_{r+1}^{(r)}(a_{1},\dots,a_{r+1})\big)=o\big(n^{r+1-1/(a_{1}\cdots a_{r})}\big),$
where $K_{r+1}^{(r)}(a_{1},\dots,a_{r+1})$ denotes the complete $(r+1)$-partite $r$-graph with $a_{1},\dots,a_{r+1}$ vertices in its parts.
Motivated by the work of Alon and Shikhelman [1], the goal of this paper is to provide a brief survey of generalized Turán problems for hypergraphs and to systematically investigate the generalized Turán problem for expansions.
2 Main results
Given $r$-graphs $\mathcal{H}$} and $\mathcal{F}$}, a generalized hypergraph Turán problem is called degenerate case if the order of magnitude of the extremal function $\mathrm{ex}_{r}(n,{\mathcal{H}},{\mathcal{F}})$ is strictly smaller than the maximum possible, namely $\mathrm{ex}_{r}(n,{\mathcal{H}},{\mathcal{F}})=o(n^{|V({\mathcal{H}})|})$. Otherwise, it is called non-degenerate case, where $\mathrm{ex}_{r}(n,{\mathcal{H}},{\mathcal{F}})=\Theta(n^{|V({\mathcal{H}})|})$. In the case ${\mathcal{H}}=K_{r}^{(r)}$, $\mathrm{ex}_{r}(n,K_{r}^{(r)},F^{r})={\rm ex}_{r}(n,F^{r})=\Theta(n^{r})$ if and only if $\chi(F)>r$, where $\chi(F)$ denotes the chromatic number of the graph $F$ [28]. In this work, we concentrate on the case where $\mathcal{H}=K_{s}^{(r)}$ (the complete $r$-graph on $s$ vertices).
2.1 The non-degenerate case
We first consider the non-degenerate case, for which we show that generalized Turán problems for expansions of graphs $F$ are non-degenerate if and only if $\chi(F)>s$.
Proposition 2.1.
Let $s\geq r$ be integers and $n$ be sufficiently large. Suppose $F$ is a graph. Then ${\rm ex}_{r}(n,K_{s}^{(r)},F^{r})=\Theta(n^{s})$ if and only if $\chi(F)>s$.
Our first main result in the non-degenerate case is an exact determination of the generalized Turán number for expansions of complete graphs, generalizing the result of Pikhurko (see Theorem 1.1).
Theorem 2.2.
Let integers $\ell\geq s\geq r\geq 3$. Then for sufficiently large $n$,
${\rm{ex}}_{r}(n,K_{s}^{(r)},K_{\ell+1}^{r})={\mathcal{N}}(K_{s}^{(r)},{\mathcal{T}}_{r}(n,\ell)).$
Mubayi [43] considered the Turán problem for the following family of $r$-graphs. For $t>r$, let ${\mathcal{K}}_{t}^{r}$ denote the family of all $r$-graphs $\mathcal{F}$} with at most $\binom{t}{2}$ hyperedges for which there exists a set $S$ of $t$ vertices, such that every pair $u,v\in S$ is contained in some hyperedge of $\mathcal{F}$}. When $r=2$, the family ${\mathcal{K}}_{t}^{r}$ reduces to $K_{t}$, however when $r>2$, it contains more than one $r$-graph. Let $\mathrm{ex}_{r}(n,K_{s}^{(r)},{\mathcal{K}}_{t}^{r})$ denote the maximum number of copies of $K_{s}^{(r)}$ in an $n$-vertex $r$-graph that contains no subhypergraph isomorphic to any member of ${\mathcal{K}}_{t}^{r}$. Mubayi [43] showed that $\mathrm{ex}_{r}(n,K_{r}^{(r)},{\mathcal{K}}_{\ell+1}^{r})=t_{r}(n,\ell)$, with the unique extremal $r$-graph being ${\mathcal{T}}_{r}(n,\ell)$. Since $K_{\ell+1}^{r}\in{\mathcal{K}}_{\ell+1}^{r}$ and ${\mathcal{T}}_{r}(n,\ell)$ is ${\mathcal{K}}_{\ell+1}^{r}$-free, applying Theorem 2.2 yields the following result.
Corollary 2.3.
Let integers $\ell\geq s\geq r\geq 3$. Then for sufficiently large $n$,
$\mathrm{ex}_{r}(n,K_{s}^{(r)},{\mathcal{K}}_{\ell+1}^{r})={\mathcal{N}}(K_{s}^{(r)},{\mathcal{T}}_{r}(n,\ell)).$
Since $K^{r}_{\ell+1}$ and ${\mathcal{K}}^{r}_{\ell+1}$ are natural generalizations of the complete graph $K_{\ell+1}$, Theorem 2.2 and Corollary 2.3 may be regarded as hypergraph extensions of the results of Zykov [53] and Erdős [14].
Our second main result establishs the asymptotics of the generalized Turán number for expansions of the vertex-disjoint union of complete graphs. Recall that $S_{n,t}^{r}(n-t,\ell)$ is obtained from ${\mathcal{T}}_{r}(n-t,\ell)$ by adding a set $U$ of $t$ new vertices and adding all $r$-sets that contain at least one vertex from $U$ as hyperedges.
Theorem 2.4.
Let integers $k\geq 1$ and $\ell\geq s\geq r\geq 3$. Then for sufficiently large $n$,
${\rm{ex}}_{r}(n,K_{s}^{(r)},kK_{\ell+1}^{r})=(1+o(1)){\mathcal{N}}\big(K_{s}^{(r)},S_{n,k-1}^{r}(n-k+1,\ell)\big).$
We remark that the condition $\ell\geq s$ in Theorems 2.2 and 2.4 is necessary. Indeed, if $\ell<s$, then these two problems become degenerate, as follows directly from Proposition 2.1.
2.2 The degenerate case
Let us turn to the degenerate case. We focus on generalized Turán problems for expansions of several classes of forests, and determine the asymptotics of generalized Turán numbers for these classes of hypergraph forests. For non-negative integers $a$ and $b$, let $\binom{a}{b}$ denote the binomial coefficient. In particular, we define $\binom{a}{b}=0$ if $a<b$, and $\binom{0}{0}=1$.
Our first two main results in this subsection concern expansions of star forests. First we establish an upper bound for ${\rm{ex}}_{r}(n,K_{s}^{(r)},S_{\ell}^{r})$, which generalizes the Duke-Erdős theorem (see Theorem 1.3).
Theorem 2.5.
Fix integers $s\geq r\geq 3$ and $\ell\geq 2$. Then there exists a constant $c(s,r,\ell)$ such that for sufficiently large $n$,
${\rm{ex}}_{r}(n,K_{s}^{(r)},S_{\ell}^{r})\leq c(s,r,\ell)n^{r-2}.$
Moreover, if $s\leq\ell+r-2$, then $\mathrm{ex}_{r}(n,K_{s}^{(r)},S_{\ell}^{r})=\Theta(n^{r-2})$.
In the general case, we determine the asymptotics of the generalized Turán number for expansions of star forests. Given an $r$-graph $\mathcal{H}$}, let ${\mathcal{F}}_{r}(n,{\mathcal{H}})$ denote the family of all $\mathcal{H}$}-free $r$-graphs on $n$ vertices.
Theorem 2.6.
Let $r\geq 3$, $k\geq 1$ and $\ell_{1}\geq\cdots\geq\ell_{k}\geq 2$ be integers and $n$ be sufficiently large.
(i) If $s\leq k+r-2$, then
${\rm{ex}}_{r}(n,K_{s}^{(r)},S^{r}_{\ell_{1}}\cup\cdots\cup S^{r}_{\ell_{k}})=\sum_{i=0}^{r-1}\binom{k-1}{s-i}\binom{n-k+1}{i}+\max_{{\mathcal{G}}\in{\mathcal{F}}_{r}(n-k+1,S^{r}_{\ell_{k}})}\sum_{i=r}^{s}\binom{k-1}{s-i}{\mathcal{N}}(K_{i}^{(r)},{\mathcal{G}}).$
(ii) If $s\geq k+r-1$, then
${\rm{ex}}_{r}(n,K_{s}^{(r)},S^{r}_{\ell_{1}}\cup\cdots\cup S^{r}_{\ell_{k}})=O(n^{r-2}).$
Moreover, if $k+r-1\leq s\leq\sum_{i=1}^{k}(\ell_{i}+1)+r-3$, then ${\rm{ex}}_{r}(n,K_{s}^{(r)},S^{r}_{\ell_{1}}\cup\cdots\cup S^{r}_{\ell_{k}})=\Theta(n^{r-2})$.
Theorem 2.6 strengthens the result of Khormali and Palmer [36] concerning star forests (see Theorem 1.6). Indeed, when $s=r$, we have $\sum_{i=0}^{r-1}\binom{k-1}{s-i}\binom{n-k+1}{i}=\binom{n}{r}-\binom{n-k+1}{r}$ and $\max_{{\mathcal{G}}\in{\mathcal{F}}_{r}(n-k+1,S^{r}_{\ell_{k}})}\sum_{i=r}^{s}\binom{k-1}{s-i}{\mathcal{N}}(K_{i}^{(r)},{\mathcal{G}})={\rm{ex}}_{r}(n-k+1,S^{r}_{\ell_{k}})$. Therefore, we have the following Turán result for expansions of star forests.
Corollary 2.7.
Let $r\geq 3$, $k\geq 1$ and $\ell_{1}\geq\cdots\geq\ell_{k}\geq 2$ be integers and $n$ be sufficiently large. Then
${\rm{ex}}_{r}(n,K_{r}^{(r)},S^{r}_{\ell_{1}}\cup\cdots\cup S^{r}_{\ell_{k}})=\binom{n}{r}-\binom{n-k+1}{r}+{\rm{ex}}_{r}(n-k+1,S^{r}_{\ell_{k}}).$
Our third main result establishes the asymptotics of generalized Turán numbers for expansions of linear forests and the vertex-disjoint union of cycles. We denoted by $C_{\ell}$ a cycle with $\ell$ edges.
Theorem 2.8.
Let $r\geq 3$, $k\geq 1$ and $\ell_{1},\dots,\ell_{k}\geq 5$ be integers and $t=\sum_{i=1}^{k}\lfloor\frac{\ell_{i}+1}{2}\rfloor-1$. Suppse $Q_{i}\in\{P_{\ell_{i}}^{r},C_{\ell_{i}}^{r}\}$, $Q_{1}\cup\cdots\cup Q_{k}\in\{P_{\ell_{1}}^{r}\cup\cdots\cup P_{\ell_{k}}^{r},C_{\ell_{1}}^{r}\cup\cdots\cup C_{\ell_{k}}^{r}\}$ and $n$ is sufficiently large.
(i) If $s\leq t+r-1$, then
${\rm{ex}}_{r}(n,K_{s}^{(r)},Q_{1}\cup\cdots\cup Q_{k})=(1+o(1)){\mathcal{N}}(K_{s}^{(r)},S_{n,t}^{r})=\binom{t}{s-r+1}\frac{n^{r-1}}{(r-1)!}+o(n^{r-1}).$
(ii) If $s\geq t+r$, then
${\rm{ex}}_{r}(n,K_{s}^{(r)},Q_{1}\cup\cdots\cup Q_{k})=o(n^{r-1}).$
In the case where $\ell_{1}=\cdots=\ell_{k}$, Theorem 2.8 immediately yields the following corollary.
Corollary 2.9.
Let $r\geq 3$, $k\geq 1$ and $\ell\geq 5$ be integers and $\ell^{\prime}=\lfloor\frac{\ell+1}{2}\rfloor$. Suppose $Q\in\{P_{\ell}^{r},C_{\ell}^{r}\}$ and $n$ is sufficiently large.
(i) If $s\leq k\ell^{\prime}+r-2$, then ${\rm{ex}}_{r}(n,K_{s}^{(r)},kQ)=(1+o(1)){\mathcal{N}}(K_{s}^{(r)},S_{n,k\ell^{\prime}-1}^{r})=\binom{k\ell^{\prime}-1}{s-r+1}\frac{n^{r-1}}{(r-1)!}+o(n^{r-1})$.
(ii) If $s\geq k\ell^{\prime}+r-1$, then ${\rm{ex}}_{r}(n,K_{s}^{(r)},kQ)=o(n^{r-1})$.
Moreover, we determine the asymptotics of generalized Turán numbers for two classes of expansions: those of star-path forests and those of the vertex-disjoint union of cycles and stars, as follows.
Theorem 2.10.
Let integers $r\geq 3$, $k\geq 1$, $p\geq 0$, $\ell_{1},\dots,\ell_{k}\geq 5$, $t_{1},\dots,t_{p}\geq 2$ if $p\geq 1$, and $t=\sum_{i=1}^{k}\lfloor\frac{\ell_{i}+1}{2}\rfloor+p-1$. Suppose $Q_{i}\in\{P_{\ell_{i}}^{r},C_{\ell_{i}}^{r}\}$, $Q_{1}\cup\cdots\cup Q_{k}\in\{P_{\ell_{1}}^{r}\cup\cdots\cup P_{\ell_{k}}^{r},C_{\ell_{1}}^{r}\cup\cdots\cup C_{\ell_{k}}^{r}\}$ and $n$ is sufficiently large.
(i) If $s\leq t+r-1$, then
${\rm{ex}}_{r}(n,K_{s}^{(r)},Q_{1}\cup\cdots\cup Q_{k}\cup S^{r}_{t_{1}}\cup\cdots\cup S^{r}_{t_{p}})=(1+o(1)){\mathcal{N}}(K_{s}^{(r)},S_{n,t}^{r})=\binom{t}{s-r+1}\frac{n^{r-1}}{(r-1)!}+o(n^{r-1}).$
(ii) If $s\geq t+r$, then
${\rm{ex}}_{r}(n,K_{s}^{(r)},Q_{1}\cup\cdots\cup Q_{k}\cup S^{r}_{t_{1}}\cup\cdots\cup S^{r}_{t_{p}})=o(n^{r-1}).$
Note that in Theorems 2.8 and 2.10, the condition $\ell_{1},\dots,\ell_{k}\geq 5$ can be relaxed to $\ell_{1},\dots,\ell_{k}\geq 4$ when $Q_{1}\cup\cdots\cup Q_{k}=P_{\ell_{1}}^{r}\cup\cdots\cup P_{\ell_{k}}^{r}$.
The rest of this paper is organized as follows. In Section 3, we present some notations and definitions, and provide the proofs of Proposition 2.1, Theorems 2.2 and 2.4. In particular, we establish a general upper bound on generalized Turán numbers for expansions (see Lemma 3.1) in Section 3. The proofs of Theorems 2.5 and 2.6 are presented in Section 4, and the proofs of Theorems 2.8 and 2.10 are presented in Section 5. Finally, we state some concluding remarks in Section 6, including several more general results.
3 Proofs of Proposition 2.1, Theorems 2.2 and 2.4
Let ${\mathcal{H}}=(V({\mathcal{H}}),E({\mathcal{H}}))$ be an $r$-graph. A strong $k$-coloring of $\mathcal{H}$} is a partition $(C_{1},C_{2},\ldots,C_{k})$ of $V({\mathcal{H}})$ such that the same color does not appear twice in the same hyperedge [6]. In other words,
$|e\cap C_{i}|\leq 1$
for any $e\in E({\mathcal{H}})$ and $i\in[k]$. The strong chromatic number of $\mathcal{H}$}, denoted by $\chi({\mathcal{H}})$, is the smallest $k$ such that $\mathcal{H}$} has a strong $k$-coloring. Clearly, $\chi(F^{r})\geq\chi(F)$ for any graph $F$.
Let $U\subseteq V({\mathcal{H}})$ be a nonempty subset. Let ${\mathcal{H}}-U$ denote the $r$-graph obtained from $\mathcal{H}$} by removing the vertices in $U$ and the hyperedges incident to them. Let ${\mathcal{H}}[U]$ denote the subhypergraph of $\mathcal{H}$} induced by $U$.
The link hypergraph of a vertex $v$ of $\mathcal{H}$} is an (r1)(r-1)-graph, defined as
[…]
The degree $d_{\mathcal{H}}(v)$ of $v$ is defined as the number of hyperedges in ${\mathcal{L}}_{v}$, i.e., the number of hyperedges in $\mathcal{H}$} that contain $v$.
Recall that $K_{s}^{(r)}$ denotes the complete $r$-graph on $s$ vertices if $s\geq r$, and an $s$-element set if $s<r$. Given two $r$-graphs ${\mathcal{G}}_{1}$ and ${\mathcal{G}}_{2}$, $\mathcal{N}({\mathcal{G}}_{1},{\mathcal{G}}_{2})$ denotes the number of copies of ${\mathcal{G}}_{1}$ in ${\mathcal{G}}_{2}$. In particular, $\mathcal{N}(K_{s}^{(r)},{\mathcal{H}})$ denotes the number of copies of $K_{s}^{(r)}$ in $\mathcal{H}$} if $s\geq r$, and the collection of all $s$-subsets of $V({\mathcal{H}})$ if $s<r$. For any $v\in V({\mathcal{H}})$ and $s\geq 1$, let
${\mathcal{K}}_{v}(K_{s}^{(r)},{\mathcal{H}})=\{K_{s}^{(r)}\subseteq{\mathcal{H}}\,|\,v\in V(K_{s}^{(r)})\},$
namely, the collection of all $K_{s}^{(r)}$ in $\mathcal{H}$} containing the vertex $v$.
For $s\geq r$, the $s$-clique hypergraph of $\mathcal{H}$} is an $s$-graph with vertex set $V({\mathcal{H}})$, where an $s$-subset of $V({\mathcal{H}})$ is a hyperedge if and only if it induces a copy of $K_{s}^{(r)}$ in $\mathcal{H}$}.
Now let us start with the proof of Proposition 2.1 that we restate here for convenience.
Proposition 2.1.
Let $s\geq r$ be integers and $n$ be sufficiently large. Suppose $F$ is a graph. Then ${\rm ex}_{r}(n,K_{s}^{(r)},F^{r})=\Theta(n^{s})$ if and only if $\chi(F)>s$.
Proof..
Recall that $\chi(F^{r})\geq\chi(F)$. Observe that we may color the r2r-2 new vertices in each hyperedge of $F^{r}$ using r2r-2 distinct colors (chosen from at most $\max\{\chi(F),r\}$ available colors), all of which differ from the colors assigned to two vertices of the edge of $F$. Then $\chi(F^{r})\leq\max\{\chi(F),r\}$. Therefore,
$\chi(F)\leq\chi(F^{r})\leq\max\{\chi(F),r\}.$ (3.1)
($\Leftarrow$). Suppose $F$ is a graph with $\chi(F)>s\geq r$. Then $\chi(F^{r})=\chi(F)>s$ by (3.1). This implies that ${\mathcal{T}}_{r}(n,s)$ is $F^{r}$-free as $\chi({\mathcal{T}}_{r}(n,s))=s$, and thus ${\rm ex}_{r}(n,K_{s}^{(r)},F^{r})\geq{\mathcal{N}}(K_{s}^{(r)},{\mathcal{T}}_{r}(n,s))=\Theta(n^{s})$. Note that ${\rm ex}_{r}(n,K_{s}^{(r)},F^{r})=O(n^{s})$. Thus, ${\rm ex}_{r}(n,K_{s}^{(r)},F^{r})=\Theta(n^{s})$.
($\Rightarrow$). Suppose now that $F$ is a graph such that ${\rm ex}_{r}(n,K_{s}^{(r)},F^{r})=\Theta(n^{s})$. Let $\mathcal{H}$} be an $F^{r}$-free $r$-graph on $n$ vertices with ${\mathcal{N}}(K_{s}^{(r)},{\mathcal{H}})={\rm ex}_{r}(n,K_{s}^{(r)},F^{r})=\Theta(n^{s})$. Observe that $\chi({\mathcal{H}})=\ell\geq s$. Indeed, if $\ell<s$, then any $s$-subset of $V({\mathcal{H}})$ contains two vertices of the same color. This implies that no $s$-subset of $V({\mathcal{H}})$ induces a copy of $K_{s}^{(r)}$. Thus, ${\mathcal{N}}(K_{s}^{(r)},{\mathcal{H}})=0$, which is a contradiction. We claim that among the $\ell$ parts of $\mathcal{H}$}, at least $s$ of them have size $\Theta(n)$. Otherwise, ${\mathcal{N}}(K_{s}^{(r)},{\mathcal{H}})=o(n^{s})$, which is a contradiction. Then $\mathcal{H}$} contains any $s$-partite $r$-graph, as $n$ is sufficiently large. This implies that $\chi(F^{r})>s$, since $\mathcal{H}$} is $F^{r}$-free. If $\chi(F)\leq s$, then $\chi(F^{r})\leq\max\{\chi(F),r\}\leq s$ by (3.1), which contradicts $\chi(F^{r})>s$. Thus, $\chi(F)>s$, completing the proof. ∎
Before proceeding to the proofs of Theorems 2.2 and 2.4, we present a lemma that establishes a general upper bound on ${\rm{ex}}_{r}(n,K_{s}^{(r)},F^{r})$ for any graph $F$.
Lemma 3.1.
Let $s\geq r$ be integers and $F$ be an arbitrary graph. Suppose $\mathcal{H}$} is an $F^{r}$-free $r$-graph on $n$ vertices. Then ${\mathcal{N}}(K_{s}^{(r)},{\mathcal{H}})\leq\mathrm{ex}_{s}(n,F^{s})$. In particular, ${\rm{ex}}_{r}(n,K_{s}^{(r)},F^{r})\leq\mathrm{ex}_{s}(n,F^{s})$.
Note that ${\mathcal{N}}(K_{s}^{(r)},{\mathcal{T}}_{r}(n,\ell))=t_{s}(n,\ell)$. Hence, Theorems 1.1 and 2.2 imply that the upper bound in Lemma 3.1 is sharp for the case when $F=K_{\ell+1}$ and $\ell\geq s$.
Proof.
Denote by ${\mathcal{H}}^{s}$ the $s$-clique hypergraph of $\mathcal{H}$}. We claim that ${\mathcal{H}}^{s}$ is $F^{s}$-free. Otherwise, assume that ${\mathcal{H}}^{s}$ contains a copy of $F^{s}$, denoted by $K$. By the definition of the $s$-clique hypergraph, every $r$-subset of each hyperedge of ${\mathcal{H}}^{s}$ is a hyperedge of $\mathcal{H}$}. Then every $r$-subset of each hyperedge of $K$ is also a hyperedge of $\mathcal{H}$}. Now we delete srs-r vertices of degree one from each hyperedge of $K$, which yields a copy of $F^{r}$, denoted by $K^{\prime}$. Since $K^{\prime}$ consists of $r$-subsets of hyperedges of $K$, $K^{\prime}$ is contained in $\mathcal{H}$}, which contradicts the fact that $\mathcal{H}$} is $F^{r}$-free. Thus, ${\mathcal{N}}(K_{s}^{(r)},{\mathcal{H}})=e({\mathcal{H}}^{s})\leq\mathrm{ex}_{s}(n,F^{s})$. ∎
We now present the proof of Theorem 2.2. Recall that it states that ${\rm{ex}}_{r}(n,K_{s}^{(r)},K_{\ell+1}^{r})={\mathcal{N}}(K_{s}^{(r)},{\mathcal{T}}_{r}(n,\ell))$ for $\ell\geq s\geq r\geq 3$ and sufficiently large $n$. The lower bound is trivially achieved by ${\mathcal{T}}_{r}(n,\ell)$. For the upper bound, combining Lemma 3.1 and Theorem 1.1, we obtain ${\rm{ex}}_{r}(n,K_{s}^{(r)},K_{\ell+1}^{r})\leq\mathrm{ex}_{s}(n,K_{\ell+1}^{s})=t_{s}(n,\ell)={\mathcal{N}}(K_{s}^{(r)},{\mathcal{T}}_{r}(n,\ell))$.
Recall that $S_{n,t}^{r}(n-t,\ell)$ is obtained from ${\mathcal{T}}_{r}(n-t,\ell)$ by adding a set $U$ of $t$ new vertices and adding all $r$-sets that contain at least one vertex from $U$ as hyperedges. We now begin with the proof of Theorem 2.4 that we restate here for convenience.
Theorem 2.4.
Let integers $k\geq 1$ and $\ell\geq s\geq r\geq 3$. Then for sufficiently large $n$,
${\rm{ex}}_{r}(n,K_{s}^{(r)},kK_{\ell+1}^{r})=(1+o(1)){\mathcal{N}}\big(K_{s}^{(r)},S_{n,k-1}^{r}(n-k+1,\ell)\big).$
Proof..
Note that ${\mathcal{T}}_{r}(n-k+1,\ell)$ is $K_{\ell+1}^{r}$-free. Hence, $S_{n,k-1}^{r}(n-k+1,\ell)$ is $kK_{\ell+1}^{r}$-free. Then ${\rm{ex}}_{r}(n,K_{s}^{(r)},kK_{\ell+1}^{r})\geq{\mathcal{N}}\big(K_{s}^{(r)},S_{n,k-1}^{r}(n-k+1,\ell)\big)$. We now consider the upper bound. Let $\mathcal{H}$} be a $kK_{\ell+1}^{r}$-free $r$-graph on $n$ vertices. Since $\ell\geq s\geq r\geq 3$, by Lemma 3.1 and Theorem 1.2, we have
${\mathcal{N}}(K_{s}^{(r)},{\mathcal{H}})\leq\mathrm{ex}_{s}(n,kK_{\ell+1}^{s})=e\big(S_{n,k-1}^{s}(n-k+1,\ell)\big)=\sum_{i=0}^{s-1}\binom{k-1}{s-i}\binom{n-k+1}{i}+t_{s}(n-k+1,\ell).$
Note that ${\mathcal{N}}\big(K_{s}^{(r)},{\mathcal{T}}_{r}(n-k+1,\ell)\big)=t_{s}(n-k+1,\ell)$. Then
${\mathcal{N}}\big(K_{s}^{(r)},S_{n,k-1}^{r}(n-k+1,\ell)\big)$ $=$ $\sum_{i=0}^{s}\binom{k-1}{s-i}{\mathcal{N}}\big(K_{i}^{(r)},{\mathcal{T}}_{r}(n-k+1,\ell)\big)$
$=$ $\sum_{i=0}^{r-1}\binom{k-1}{s-i}\binom{n-k+1}{i}+\sum_{i=r}^{s-1}\binom{k-1}{s-i}{\mathcal{N}}\big(K_{i}^{(r)},{\mathcal{T}}_{r}(n-k+1,\ell)\big)$
$+t_{s}(n-k+1,\ell).$
Thus,
${\mathcal{N}}(K_{s}^{(r)},{\mathcal{H}})\leq\sum_{i=0}^{s-1}\binom{k-1}{s-i}\binom{n-k+1}{i}+t_{s}(n-k+1,\ell)=(1+o(1)){\mathcal{N}}\big(K_{s}^{(r)},S_{n,k-1}^{r}(n-k+1,\ell)\big).$
This completes the proof. ∎
4 Proofs of Theorems 2.5 and 2.6
Now let us start with the proof of Theorem 2.5 that we restate here for convenience.
Theorem 2.5.
Fix integers $s\geq r\geq 3$ and $\ell\geq 2$. Then there exists a constant $c(s,r,\ell)$ such that for sufficiently large $n$,
${\rm{ex}}_{r}(n,K_{s}^{(r)},S_{\ell}^{r})\leq c(s,r,\ell)n^{r-2}.$
Moreover, if $s\leq\ell+r-2$, then $\mathrm{ex}_{r}(n,K_{s}^{(r)},S_{\ell}^{r})=\Theta(n^{r-2})$.
Proof..
Let $\mathcal{H}$} be an $r$-graph on $n$ vertices with ${\mathcal{N}}(K_{s}^{(r)},{\mathcal{H}})>c(s,r,\ell)n^{r-2}$ for some sufficiently large constant $c(s,r,\ell)$. We will show that $\mathcal{H}$} contains a copy of $S_{\ell}^{r}$.
Claim 4.1.
$|{\mathcal{K}}_{v}(K_{s}^{(r)},{\mathcal{H}})|=O(n^{r-2})$ for any $v\in V({\mathcal{H}})$.
Proof of Claim.
Suppose to the contrary that there exists a vertex $u\in V({\mathcal{H}})$ such that $|{\mathcal{K}}_{u}(K_{s}^{(r)},{\mathcal{H}})|=\omega(n^{r-2})$. This implies that there are $\omega(n^{r-2})$ copies of $K_{s-1}^{(r)}$ such that each $K_{s-1}^{(r)}$ forms a $K_{s}^{(r)}$ together with $u$. Denote by $X_{s-1}$ the collection of such $K_{s-1}^{(r)}$. Then $|X_{s-1}|=\omega(n^{r-2})$.
Let ${\mathcal{S}}^{r-1}=\big\{K_{r-1}^{(r)}\ \big|\ K_{r-1}^{(r)}\subseteq V(K_{s-1}^{(r)})\ \text{and}\ K_{s-1}^{(r)}\in X_{s-1}\big\}$. Consider the (r1)(r-1)-graph ${\mathcal{H}}^{r-1}=\Big(\bigcup_{K_{s-1}^{(r)}\in X_{s-1}}V(K_{s-1}^{(r)}),{\mathcal{S}}^{r-1}\Big)$. We claim that ${\mathcal{H}}^{r-1}$ is $M_{\ell}^{r-1}$-free. Otherwise, assume that ${\mathcal{H}}^{r-1}$ contains a copy of $M_{\ell}^{r-1}$. By the definitions of $X_{s-1}$ and ${\mathcal{S}}^{r-1}$, the vertex $u$ forms a hyperedge of $\mathcal{H}$} together with any (r1)(r-1)-edge of ${\mathcal{H}}^{r-1}$. Thus, ${\mathcal{H}}[\{u\}\cup V(M_{\ell}^{r-1})]$ contains a copy of $S_{\ell}^{r}$ with $u$ as its center (i.e., the unique vertex of degree not equal to one), and we are done.
Since ${\mathcal{H}}^{r-1}$ is $M_{\ell}^{r-1}$-free, by Theorem 1.8, we have ${\mathcal{N}}(K_{s-1}^{(r-1)},{\mathcal{H}}^{r-1})=O(n^{r-2})$. By the definitions of ${\mathcal{S}}^{r-1}$ and ${\mathcal{H}}^{r-1}$, it follows that $|X_{s-1}|\leq{\mathcal{N}}(K_{s-1}^{(r-1)},{\mathcal{H}}^{r-1})$, since each $K_{s-1}^{(r)}\in X_{s-1}$ corresponds to a copy of $K_{s-1}^{(r-1)}$ in ${\mathcal{H}}^{r-1}$. Thus, $|X_{s-1}|=O(n^{r-2})$, which contradicts $|X_{s-1}|=\omega(n^{r-2})$. This completes the proof of Claim 4.1. ∎
[…]
Now we delete the vertex set $S_{s-r+1}\cup S_{r-3}\cup\{u\}$ from $\mathcal{H}$}. Consequently, at most $(s-1)\lambda n^{r-3}$ copies of $K_{s}^{(r)}$ in $\mathcal{H}$} are removed, and at least $\lambda/\binom{(r-1)(\ell-1)}{s-r+1}$ copies of $K_{s}^{(r)}$ in ${\mathcal{K}}_{u}(K_{s}^{(r)},{\mathcal{H}})$ containing $S_{s-r+1}\cup S_{r-3}\cup\{u\}$ are deleted. Since $c(s,r,\ell)$ is sufficiently large, we may repeat this procedure (more than $\ell$ times) until $c(s,r,\ell)n^{r-2}/2$ copies of $K_{s}^{(r)}$ in $\mathcal{H}$} have been deleted. More precisely, we repeat this procedure at least $c(s,r,\ell)n^{r-2}/(2(s-1)\lambda n^{r-3})=c(s,r,\ell)n/(2(s-1)\lambda)$ times. In the process, we delete at least $c(s,r,\ell)n/(2(s-1)\lambda)$ distinct (s1)(s-1)-sets, and thus remove at least $\frac{c(s,r,\ell)n}{2(s-1)\lambda}\cdot\lambda\cdot\binom{(r-1)(\ell-1)}{s-r+1}^{-1}=\frac{c(s,r,\ell)n}{2(s-1)}\cdot\binom{(r-1)(\ell-1)}{s-r+1}^{-1}$ copies of $K_{s}^{(r)}$ in $\mathcal{H}$} containing some deleted (s1)(s-1)-set. Without loss of generality, let $Y_{s-1}$ be the collection of all deleted (s1)(s-1)-sets. Then $|Y_{s-1}|\geq\frac{c(s,r,\ell)n}{2(s-1)\lambda}>\ell$.
We claim that there exists a vertex $w\in V({\mathcal{H}})$ such that $w$ forms a copy of $K_{s}^{(r)}$ in $\mathcal{H}$} with each of $\ell$ distinct (s1)(s-1)-sets in $Y_{s-1}$. Otherwise, every vertex in $V({\mathcal{H}})$ forms a copy of $K_{s}^{(r)}$ in $\mathcal{H}$} with at most $\ell-1$ distinct (s1)(s-1)-sets in $Y_{s-1}$. This implies that we can only remove at most $(\ell-1)n$ copies of $K_{s}^{(r)}$ in $\mathcal{H}$} containing some (s1)(s-1)-set in $Y_{s-1}$. However, it follows that $\frac{c(s,r,\ell)}{2(s-1)}\cdot\binom{(r-1)(\ell-1)}{s-r+1}^{-1}>\ell-1$ as $c(s,r,\ell)$ is sufficiently large, which is a contradiction. Thus, by selecting one (r1)(r-1)-subset from each of $\ell$ distinct (s1)(s-1)-sets in $Y_{s-1}$, we may obtain a copy of $S_{\ell}^{r}$ with $w$ as its center, completing the proof.
If $s\leq\ell+r-2$, then we consider the following $r$-graph ${\mathcal{H}}_{1}$: take a vertex set partition $A\cup B$ where $|A|=\ell$ and $|B|=n$-$\ell$, then add as hyperedges all $r$-sets containing at least two vertices from $A$. We claim that ${\mathcal{H}}_{1}$ is $S_{\ell}^{r}$-free. Suppose to the contrary that ${\mathcal{H}}_{1}$ contains a copy of $S_{\ell}^{r}$. Let $u\in V(S_{\ell}^{r})$ denote the center of this $S_{\ell}^{r}$. If $u\in A$, then by the definition of ${\mathcal{H}}_{1}$, $|V(S_{\ell}^{r})\cap A|\geq\ell+1$, contradicting $|A|=\ell$. If $u\in B$, then by the definition of ${\mathcal{H}}_{1}$, $|V(S_{\ell}^{r})\cap A|\geq 2\ell$, which also contradicts $|A|=\ell$. Therefore, ${\mathcal{H}}_{1}$ is $S_{\ell}^{r}$-free. Then
$\mathrm{ex}_{r}(n,K_{s}^{(r)},S_{\ell}^{r})\geq{\mathcal{N}}(K_{s}^{(r)},{\mathcal{H}}_{1})=\sum_{i=0}^{r-2}\binom{\ell}{s-r+2+i}\binom{n-\ell}{r-2-i}$.
Combining the upper bound, we have $\mathrm{ex}_{r}(n,K_{s}^{(r)},S_{\ell}^{r})=\Theta(n^{r-2})$. ∎
Recall that for an $r$-graph $\mathcal{H}$}, ${\mathcal{F}}_{r}(n,{\mathcal{H}})$ denotes the family of all $\mathcal{H}$}-free $r$-graphs on $n$ vertices. We now continue with the proof of Theorem 2.6 that we restate here for convenience.
Theorem 2.6.
Let $r\geq 3$, $k\geq 1$ and $\ell_{1}\geq\cdots\geq\ell_{k}\geq 2$ be integers and $n$ be sufficiently large.
(i) If $s\leq k+r-2$, then
${\rm{ex}}_{r}(n,K_{s}^{(r)},S^{r}_{\ell_{1}}\cup\cdots\cup S^{r}_{\ell_{k}})=\sum_{i=0}^{r-1}\binom{k-1}{s-i}\binom{n-k+1}{i}+\max_{{\mathcal{G}}\in{\mathcal{F}}_{r}(n-k+1,S^{r}_{\ell_{k}})}\sum_{i=r}^{s}\binom{k-1}{s-i}{\mathcal{N}}(K_{i}^{(r)},{\mathcal{G}}).$
(ii) If $s\geq k+r-1$, then
${\rm{ex}}_{r}(n,K_{s}^{(r)},S^{r}_{\ell_{1}}\cup\cdots\cup S^{r}_{\ell_{k}})=O(n^{r-2}).$
Moreover, if $k+r-1\leq s\leq\sum_{i=1}^{k}(\ell_{i}+1)+r-3$, then ${\rm{ex}}_{r}(n,K_{s}^{(r)},S^{r}_{\ell_{1}}\cup\cdots\cup S^{r}_{\ell_{k}})=\Theta(n^{r-2})$.
Proof..
For the lower bound in (i), we consider the following $r$-graph. Let us take any $S^{r}_{\ell_{k}}$-free $r$-graph $\mathcal{G}$} on nk+1n-k+1 vertices and a set $U$ of k1k-1 vertices, then add each hyperedge that contains at least one vertex of $U$. Since $\mathcal{G}$} is $S^{r}_{\ell_{k}}$-free and $\ell_{1}\geq\cdots\geq\ell_{k}$, any $S_{\ell_{i}}^{r}$ in the resulting $r$-graph contains at least one vertex in $U$. Thus, the resulting $r$-graph is $S^{r}_{\ell_{1}}\cup\cdots\cup S^{r}_{\ell_{k}}$-free as $|U|<k$. Clearly, the number of copies of $K_{s}^{(r)}$ in the resulting $r$-graph is $\sum_{i=0}^{r-1}\binom{k-1}{s-i}\binom{n-k+1}{i}+\sum_{i=r}^{s}\binom{k-1}{s-i}{\mathcal{N}}(K_{i}^{(r)},{\mathcal{G}})$.
We now continue with the upper bound in (i). Suppose $\mathcal{H}$} is an $r$-graph on $n$ vertices with
${\mathcal{N}}(K_{s}^{(r)},{\mathcal{H}})>\sum_{i=0}^{r-1}\binom{k-1}{s-i}\binom{n-k+1}{i}+\max_{{\mathcal{G}}\in{\mathcal{F}}_{r}(n-k+1,S^{r}_{\ell_{k}})}\sum_{i=r}^{s}\binom{k-1}{s-i}{\mathcal{N}}(K_{i}^{(r)},{\mathcal{G}}).$
We will show that $\mathcal{H}$} contains a copy of $S^{r}_{\ell_{1}}\cup\cdots\cup S^{r}_{\ell_{k}}$ by induction on $k$. If $k=1$, then ${\mathcal{N}}(K_{s}^{(r)},{\mathcal{H}})>\max_{{\mathcal{G}}\in{\mathcal{F}}_{r}(n,S^{r}_{\ell_{1}})}{\mathcal{N}}(K_{s}^{(r)},{\mathcal{G}})=\mathrm{ex}_{r}(n,K_{s}^{(r)},S^{r}_{\ell_{1}})$, and thus $\mathcal{H}$} contains a copy of $S_{\ell_{1}}^{r}$. Suppose $k\geq 2$ and the statement holds for any $k^{\prime}<k$. By Theorem 2.5, there exists a constant $c(s,r,\ell_{1})$ such that
${\rm{ex}}_{r}(n,K_{s}^{(r)},S_{\ell_{1}}^{r})\leq c(s,r,\ell_{1})n^{r-2}.$ (4.1)
Let $u\in V({\mathcal{H}})$ be a vertex such that $|{\mathcal{K}}_{u}(K_{s}^{(r)},{\mathcal{H}})|=\max_{v\in V({\mathcal{H}})}|{\mathcal{K}}_{v}(K_{s}^{(r)},{\mathcal{H}})|$. Now we consider the following two cases.
Case 1. $|{\mathcal{K}}_{u}(K_{s}^{(r)},{\mathcal{H}})|<\frac{{\mathcal{N}}(K_{s}^{(r)},{\mathcal{H}})-c(s,r,\ell_{1})n^{r-2}}{\sum_{i=2}^{k}\left(\ell_{i}(r-1)+1\right)}$.
By Theorem 2.5, for any ${\mathcal{G}}\in{\mathcal{F}}_{r}(n-k+1,S^{r}_{\ell_{k}})$ and all $r\leq i\leq s$, we have ${\mathcal{N}}(K_{i}^{(r)},{\mathcal{G}})=O(n^{r-2})$. Therefore, $\max_{{\mathcal{G}}\in{\mathcal{F}}_{r}(n-k+1,S^{r}_{\ell_{k}})}\sum_{i=r}^{s}\binom{k-1}{s-i}{\mathcal{N}}(K_{i}^{(r)},{\mathcal{G}})=O(n^{r-2})$. Since $s\leq k+r-2$ and $n$ is sufficiently large,
${\mathcal{N}}(K_{s}^{(r)},{\mathcal{H}})$ $>$ $\sum_{i=0}^{r-1}\binom{k-1}{s-i}\binom{n-k+1}{i}+\max_{{\mathcal{G}}\in{\mathcal{F}}_{r}(n-k+1,S^{r}_{\ell_{k}})}\sum_{i=r}^{s}\binom{k-1}{s-i}{\mathcal{N}}(K_{i}^{(r)},{\mathcal{G}})$
$\geq$ $\binom{k-1}{s-r+1}\binom{n-k+1}{r-1}$
$>$ $\sum_{i=0}^{r-1}\binom{k-2}{s-i}\binom{n-k+2}{i}+\max_{{\mathcal{G}}\in{\mathcal{F}}_{r}(n-k+1,S^{r}_{\ell_{k}})}\sum_{i=r}^{s}\binom{k-1}{s-i}{\mathcal{N}}(K_{i}^{(r)},{\mathcal{G}}).$
By induction hypothesis, $\mathcal{H}$} contains a copy of $S^{r}_{\ell_{2}}\cup\cdots\cup S^{r}_{\ell_{k}}$, denoted by $S$. Therefore, $|V(S)|=\sum_{i=2}^{k}\left(\ell_{i}(r-1)+1\right)$. Then
${\mathcal{N}}(K_{s}^{(r)},{\mathcal{H}}-V(S))$ $>$ ${\mathcal{N}}(K_{s}^{(r)},{\mathcal{H}})-|V(S)|\cdot\frac{{\mathcal{N}}(K_{s}^{(r)},{\mathcal{H}})-c(s,r,\ell_{1})n^{r-2}}{\sum_{i=2}^{k}\left(\ell_{i}(r-1)+1\right)}$
$=$ $c(s,r,\ell_{1})n^{r-2},$
which implies ${\mathcal{H}}-V(S)$ contains a copy of $S_{\ell_{1}}^{r}$ by (4.1). Thus, $\mathcal{H}$} contains a copy of $S^{r}_{\ell_{1}}\cup\cdots\cup S^{r}_{\ell_{k}}$.
Case 2. $|{\mathcal{K}}_{u}(K_{s}^{(r)},{\mathcal{H}})|\geq\frac{{\mathcal{N}}(K_{s}^{(r)},{\mathcal{H}})-c(s,r,\ell_{1})n^{r-2}}{\sum_{i=2}^{k}\left(\ell_{i}(r-1)+1\right)}$.
Recall that ${\mathcal{N}}(K_{s}^{(r)},{\mathcal{H}})>\binom{k-1}{s-r+1}\binom{n-k+1}{r-1}$. Therefore, $|{\mathcal{K}}_{u}(K_{s}^{(r)},{\mathcal{H}})|=\Omega(n^{r-1})$. Set
${\mathcal{K}}^{s-1}=\big\{K_{s}^{(r)}-\{u\}\ \big|\ K_{s}^{(r)}\in{\mathcal{K}}_{u}(K_{s}^{(r)},{\mathcal{H}})\big\}.$
Then $|{\mathcal{K}}^{s-1}|=|{\mathcal{K}}_{u}(K_{s}^{(r)},{\mathcal{H}})|=\Omega(n^{r-1})$.
Claim 4.2.
${\mathcal{H}}-\{u\}$ is $S^{r}_{\ell_{2}}\cup\cdots\cup S^{r}_{\ell_{k}}$-free.
Proof of Claim.
Suppose to the contrary that ${\mathcal{H}}-\{u\}$ contains a copy of $S^{r}_{\ell_{2}}\cup\cdots\cup S^{r}_{\ell_{k}}$, denoted by $S^{\prime}$. Therefore, $|V(S^{\prime})|=\sum_{i=2}^{k}\left(\ell_{i}(r-1)+1\right)$. Observe that
$\big|\big\{K_{s-1}^{(r)}\in{\mathcal{K}}^{s-1}\,\big|\,|V(K_{s-1}^{(r)})\cap V(S^{\prime})|\geq s-r+1\big\}\big|=O(n^{r-2}).$
Then there are at least $|{\mathcal{K}}^{s-1}|-O(n^{r-2})=\Omega(n^{r-1})$ copies of $K_{s-1}^{(r)}$ in ${\mathcal{K}}^{s-1}$ such that $|V(K_{s-1}^{(r)})\backslash V(S^{\prime})|\geq r-1$. For $r-1\leq i\leq s-1$, set
$X_{i}=\big\{K_{s-1}^{(r)}\in{\mathcal{K}}^{s-1}\,\big|\,|V(K_{s-1}^{(r)})\backslash V(S^{\prime})|=i\big\}.$
Then $\sum_{i=r-1}^{s-1}|X_{i}|=\Omega(n^{r-1})$. Now we define the following collections. Let
$X^{\prime}_{i}=\big\{K_{s-1}^{(r)}-V(S^{\prime})\,\big|\,K_{s-1}^{(r)}\in X_{i}\big\}$
for $r-1\leq i\leq s-1$. Note that $X^{\prime}_{i}$ is a collection of (r1)(r-1)-sets in $V({\mathcal{H}})\backslash(\{u\}\cup V(S^{\prime}))$ when i=r1i=r-1, and $X^{\prime}_{i}$ is a collection of $K_{i}^{(r)}$ in $V({\mathcal{H}})\backslash(\{u\}\cup V(S^{\prime}))$ when $r-1<i\leq s-1$. By the definitions of $X_{i}$ and $X^{\prime}_{i}$, we have $|X^{\prime}_{i}|\geq|X_{i}|/\binom{|V(S^{\prime})|}{s-i-1}$ for each $r-1\leq i\leq s-1$. Thus,
$\sum_{i=r-1}^{s-1}|X^{\prime}_{i}|$ $\geq$ $\sum_{i=r-1}^{s-1}|X_{i}|\cdot\binom{|V(S^{\prime})|}{s-i-1}^{-1}$
$\geq$ $\min_{r-1\leq i\leq s-1}\binom{|V(S^{\prime})|}{s-i-1}^{-1}\cdot\left(|X_{r-1}|+|X_{r}|+\cdots+|X_{s-1}|\right)$
$=$ $\Omega(n^{r-1}).$
By the pigeonhole principle, $|X^{\prime}_{i}|=\Omega(n^{r-1})$ for some $r-1\leq i\leq s-1$. Note that $i$ is fixed in what follows.
Let ${\mathcal{S}}^{r-1}=\big\{K_{r-1}^{(r)}\ \big|\ K_{r-1}^{(r)}\subseteq V(K_{j}^{(r)})\ \text{and}\ K_{j}^{(r)}\in X^{\prime}_{j},\,r-1\leq j\leq s-1\big\}$. Consider the (r1)(r-1)-graph ${\mathcal{H}}^{r-1}=\left(V({\mathcal{H}})\backslash(\{u\}\cup V(S^{\prime})),{\mathcal{S}}^{r-1}\right)$. We claim that ${\mathcal{H}}^{r-1}$ is $M_{\ell_{1}}^{r-1}$-free. Otherwise, assume that ${\mathcal{H}}^{r-1}$ contains a copy of $M_{\ell_{1}}^{r-1}$. Note that each $K_{j}^{(r)}\in X^{\prime}_{j}$ forms a $K_{j+1}^{(r)}$ together with the vertex $u$. This implies that ${\mathcal{H}}[\{u\}\cup V(M_{\ell_{1}}^{r-1})]$ contains a copy of $S_{\ell_{1}}^{r}$ with $u$ as its center. Thus, $S_{\ell_{1}}^{r}\cup S^{\prime}$ forms a copy of $S^{r}_{\ell_{1}}\cup\cdots\cup S^{r}_{\ell_{k}}$ in $\mathcal{H}$}, and we are done.
Since ${\mathcal{H}}^{r-1}$ is $M_{\ell_{1}}^{r-1}$-free, by Theorem 1.8, we have ${\mathcal{N}}(K_{i}^{(r-1)},{\mathcal{H}}^{r-1})=O(n^{r-2})$. By the definitions of ${\mathcal{S}}^{r-1}$ and ${\mathcal{H}}^{r-1}$, it follows that $|X^{\prime}_{i}|\leq{\mathcal{N}}(K_{i}^{(r-1)},{\mathcal{H}}^{r-1})$, since each $K_{i}^{(r)}\in X^{\prime}_{i}$ corresponds to a copy of $K_{i}^{(r-1)}$ in ${\mathcal{H}}^{r-1}$. Thus, $|X^{\prime}_{i}|=O(n^{r-2})$, which contradicts $|X^{\prime}_{i}|=\Omega(n^{r-1})$. This completes the proof of Claim 4.2. ∎
By Claim 4.2 and the induction hypothesis, we have
${\mathcal{N}}(K_{s-1}^{(r)},{\mathcal{H}}-\{u\})$ $\leq$ $\sum_{i=0}^{r-1}\binom{k-2}{s-i-1}\binom{n-k+1}{i}+\max_{{\mathcal{G}}\in{\mathcal{F}}_{r}(n-k+1,S^{r}_{\ell_{k}})}\sum_{i=r}^{s}\binom{k-2}{s-i-1}{\mathcal{N}}(K_{i}^{(r)},{\mathcal{G}}).$
Therefore,
$|{\mathcal{K}}_{u}(K_{s}^{(r)},{\mathcal{H}})|$ $\leq$ ${\mathcal{N}}(K_{s-1}^{(r)},{\mathcal{H}}-\{u\})$
$\leq$ $\sum_{i=0}^{r-1}\binom{k-2}{s-i-1}\binom{n-k+1}{i}+\max_{{\mathcal{G}}\in{\mathcal{F}}_{r}(n-k+1,S^{r}_{\ell_{k}})}\sum_{i=r}^{s}\binom{k-2}{s-i-1}{\mathcal{N}}(K_{i}^{(r)},{\mathcal{G}}).$
Thus,
${\mathcal{N}}(K_{s}^{(r)},{\mathcal{H}}-\{u\})$ $=$ ${\mathcal{N}}(K_{s}^{(r)},{\mathcal{H}})-|{\mathcal{K}}_{u}(K_{s}^{(r)},{\mathcal{H}})|$
$>$ $\sum_{i=0}^{r-1}\binom{k-1}{s-i}\binom{n-k+1}{i}+\max_{{\mathcal{G}}\in{\mathcal{F}}_{r}(n-k+1,S^{r}_{\ell_{k}})}\sum_{i=r}^{s}\binom{k-1}{s-i}{\mathcal{N}}(K_{i}^{(r)},{\mathcal{G}})$
$-\sum_{i=0}^{r-1}\binom{k-2}{s-i-1}\binom{n-k+1}{i}-\max_{{\mathcal{G}}\in{\mathcal{F}}_{r}(n-k+1,S^{r}_{\ell_{k}})}\sum_{i=r}^{s}\binom{k-2}{s-i-1}{\mathcal{N}}(K_{i}^{(r)},{\mathcal{G}})$
$=$ $\sum_{i=0}^{r-1}\binom{k-2}{s-i}\binom{n-k+1}{i}+\max_{{\mathcal{G}}\in{\mathcal{F}}_{r}(n-k+1,S^{r}_{\ell_{k}})}\sum_{i=r}^{s}\binom{k-2}{s-i}{\mathcal{N}}(K_{i}^{(r)},{\mathcal{G}}).$
This implies that ${\mathcal{H}}-\{u\}$ contains a copy of $S^{r}_{\ell_{2}}\cup\cdots\cup S^{r}_{\ell_{k}}$ by the induction hypothesis, which contradicts Claim 4.2. This completes the proof of (i).
Now we consider the upper bound in (ii). Suppose $\mathcal{H}$} is an $r$-graph on $n$ vertices with ${\mathcal{N}}(K_{s}^{(r)},{\mathcal{H}})=\omega(n^{r-2})$. When $k=1$, the conclusion follows from Theorem 2.5. Analogous to the proof of the upper bound in (i), we may show that $\mathcal{H}$} contains a copy of $S^{r}_{\ell_{1}}\cup\cdots\cup S^{r}_{\ell_{k}}$ by induction on $k$, completing the proof.
If $k+r-1\leq s\leq\sum_{i=1}^{k}(\ell_{i}+1)+r-3$, then we consider the following $r$-graph ${\mathcal{H}}_{2}$: take a vertex set partition $A\cup B$ where $|A|=\sum_{i=1}^{k}(\ell_{i}+1)-1$ and $|B|=n-\sum_{i=1}^{k}(\ell_{i}+1)+1$, then add as hyperedges all $r$-sets containing at least two vertices from $A$. We claim that ${\mathcal{H}}_{2}$ is $S^{r}_{\ell_{1}}\cup\cdots\cup S^{r}_{\ell_{k}}$-free. Suppose to the contrary that ${\mathcal{H}}_{2}$ contains a copy of $S^{r}_{\ell_{1}}\cup\cdots\cup S^{r}_{\ell_{k}}$, denoted by $\mathcal{C}$}. Similar to the analysis in Theorem 2.5, we have $|V(S_{\ell_{i}}^{r})\cap A|\geq\min\{\ell_{i}+1,2\ell_{i}\}=\ell_{i}+1$. Then $|V({\mathcal{C}})\cap A|\geq\sum_{i=1}^{k}(\ell_{i}+1)$, which contradicts $|A|=\sum_{i=1}^{k}(\ell_{i}+1)-1$. Therefore, ${\mathcal{H}}_{2}$ is $S^{r}_{\ell_{1}}\cup\cdots\cup S^{r}_{\ell_{k}}$-free. Then
$\mathrm{ex}_{r}(n,K_{s}^{(r)},S_{\ell}^{r})\geq{\mathcal{N}}(K_{s}^{(r)},{\mathcal{H}}_{2})=\sum_{i=0}^{r-2}\binom{\sum_{i=1}^{k}(\ell_{i}+1)-1}{s-r+2+i}\binom{n-\sum_{i=1}^{k}(\ell_{i}+1)+1}{r-2-i}$.
Combining the upper bound, we have $\mathrm{ex}_{r}(n,K_{s}^{(r)},S^{r}_{\ell_{1}}\cup\cdots\cup S^{r}_{\ell_{k}})=\Theta(n^{r-2})$. ∎
5 Proofs of Theorems 2.8 and 2.10
In [51], we determined the exact Turán number for the vertex-disjoint union of expansions of paths and cycles. For the sake of completeness, a proof of this result is included in the Appendix.
Theorem 5.1 (Zhou and Yuan [51]).
Let integers $r\geq 3$, $k\geq 1$, $0\leq p\leq k$ and $\ell_{1},\dots,\ell_{k}\geq 3$, where $\ell_{i}\neq 4$ for all $i$ if $r=3$. Then for sufficiently large $n$,
${\rm{ex}}_{r}(n,P_{\ell_{1}}^{r}\cup\cdots\cup P_{\ell_{p}}^{r}\cup C_{\ell_{p+1}}^{r}\cup\cdots\cup C_{\ell_{k}}^{r})=\binom{n}{r}-\binom{n-\sum_{i=1}^{k}\big\lfloor\frac{\ell_{i}+1}{2}\big\rfloor+1}{r}+c(n;\ell_{1},\dots,\ell_{k}),$
where $c(n;\ell_{1},\dots,\ell_{k})=\binom{n-\sum_{i=1}^{k}\lfloor\frac{\ell_{i}+1}{2}\rfloor-1}{r-2}$ if all of $\ell_{1},\dots,\ell_{k}$ are even, and $c(n;\ell_{1},\dots,\ell_{k})=0$ otherwise.
Recall that $S_{n,t}^{r}$ denotes the $n$-vertex $r$-graph consisting of all hyperedges that intersect a fixed set of $t$ vertices. Now let us start with the proof of Theorem 2.8 that we restate here for convenience.
Theorem 2.8.
Let $r\geq 3$, $k\geq 1$ and $\ell_{1},\dots,\ell_{k}\geq 5$ be integers and $t=\sum_{i=1}^{k}\lfloor\frac{\ell_{i}+1}{2}\rfloor-1$. Suppose $Q_{i}\in\{P_{\ell_{i}}^{r},C_{\ell_{i}}^{r}\}$, $Q_{1}\cup\cdots\cup Q_{k}\in\{P_{\ell_{1}}^{r}\cup\cdots\cup P_{\ell_{k}}^{r},C_{\ell_{1}}^{r}\cup\cdots\cup C_{\ell_{k}}^{r}\}$ and $n$ is sufficiently large.
(i) If $s\leq t+r-1$, then
${\rm{ex}}_{r}(n,K_{s}^{(r)},Q_{1}\cup\cdots\cup Q_{k})=(1+o(1)){\mathcal{N}}(K_{s}^{(r)},S_{n,t}^{r})=\binom{t}{s-r+1}\frac{n^{r-1}}{(r-1)!}+o(n^{r-1}).$
(ii) If $s\geq t+r$, then
${\rm{ex}}_{r}(n,K_{s}^{(r)},Q_{1}\cup\cdots\cup Q_{k})=o(n^{r-1}).$
Proof..
For the lower bound, we consider the $r$-graph $S_{n,t}^{r}$. Observe that any $P_{\ell_{i}}^{r}$ or $C_{\ell_{i}}^{r}$ in $S_{n,t}^{r}$ contains at least $\big\lfloor\frac{\ell_{i}+1}{2}\big\rfloor$ vertices in the $t$-set of $S_{n,t}^{r}$. Thus, $S_{n,t}^{r}$ is $Q_{1}\cup\cdots\cup Q_{k}$-free as $t<\sum_{i=1}^{k}\lfloor\frac{\ell_{i}+1}{2}\rfloor$.
We now prove the upper bounds in (i) and (ii) together, proceeding by double induction on $s$ and $k$. For the case when $k=1$ and $s\geq r$, the conclusion follows from Theorem 1.10. For the case when $s=r$ and $k\geq 1$, the conclusion follows from Theorems 1.5 (when $Q_{1}\cup\cdots\cup Q_{k}=P_{\ell_{1}}^{r}\cup\cdots\cup P_{\ell_{k}}^{r}$) and 5.1 (when $Q_{1}\cup\cdots\cup Q_{k}=C_{\ell_{1}}^{r}\cup\cdots\cup C_{\ell_{k}}^{r}$). Now we assume that $s>r$, $k>1$ and the conclusion holds for any $s^{\prime}\leq s$ and k1k-1. Let $\mathcal{H}$} be an $r$-graph on $n$ vertices with
${\mathcal{N}}(K_{s}^{(r)},{\mathcal{H}})>\binom{t}{s-r+1}\frac{n^{r-1}}{(r-1)!}+o(n^{r-1}).$
This implies that ${\mathcal{N}}(K_{s}^{(r)},{\mathcal{H}})\sim an^{r-1}$ as $n$ is sufficiently large, where $a>\binom{t}{s-r+1}\frac{1}{(r-1)!}$ if $s\leq t+r-1$, and $a>0$ if $s\geq t+r$. Next, we will show that $\mathcal{H}$} contains a copy of $Q_{1}\cup\cdots\cup Q_{k}$. For convenience, let $\ell^{\prime}=\lfloor\frac{\ell_{1}+1}{2}\big\rfloor$. Since $\binom{t}{s-r+1}\geq\binom{\ell^{\prime}-1}{s-r+1}$ and $n$ is sufficiently large,
${\mathcal{N}}(K_{s}^{(r)},{\mathcal{H}})\sim an^{r-1}>\binom{t}{s-r+1}\frac{n^{r-1}}{(r-1)!}+o(n^{r-1})\geq\binom{\ell^{\prime}-1}{s-r+1}\frac{n^{r-1}}{(r-1)!}+o(n^{r-1}),$
which implies $\mathcal{H}$} contains a copy of $Q_{1}$ by Theorem 1.10, denoted by $F$. Therefore, $|V(F)|=\ell_{1}(r-1)+1$ if $Q_{1}=P_{\ell_{1}}^{r}$, and $|V(F)|=\ell_{1}(r-1)$ if $Q_{1}=C_{\ell_{1}}^{r}$.
We may assume that
${\mathcal{N}}\big(K_{s}^{(r)},{\mathcal{H}}-V(F)\big)\leq\binom{t-\ell^{\prime}}{s-r+1}\frac{n^{r-1}}{(r-1)!}+o(n^{r-1}).$
Otherwise, by the induction hypothesis, we may find a copy of $Q_{2}\cup\cdots\cup Q_{k}$ in ${\mathcal{H}}-V(F)$. Thus, $\mathcal{H}$} contains a copy of $Q_{1}\cup\cdots\cup Q_{k}$ and we are done.
Let ${\mathcal{K}}_{V(F)}(K_{s}^{(r)},{\mathcal{H}})=\{K_{s}^{(r)}\subseteq{\mathcal{H}}\,|\,V(K_{s}^{(r)})\cap V(F)\neq\emptyset\}$, namely, the collection of all $K_{s}^{(r)}$ in $\mathcal{H}$} incident to vertices in $V(F)$. Then
$|{\mathcal{K}}_{V(F)}(K_{s}^{(r)},{\mathcal{H}})|$ $=$ ${\mathcal{N}}(K_{s}^{(r)},{\mathcal{H}})-{\mathcal{N}}(K_{s}^{(r)},{\mathcal{H}}-V(F))$ (5.1)
$>$ $\left(\binom{t}{s-r+1}-\binom{t-\ell^{\prime}}{s-r+1}\right)\frac{n^{r-1}}{(r-1)!}+o(n^{r-1}).$
Now we estimate the upper bound for $|{\mathcal{K}}_{V(F)}(K_{s}^{(r)},{\mathcal{H}})|$. We partition ${\mathcal{K}}_{V(F)}(K_{s}^{(r)},{\mathcal{H}})$ into the following sr+2s-r+2 parts. Let
${\mathcal{K}}^{*}_{V(F)}(K_{s}^{(r)},{\mathcal{H}})=\big\{K_{s}^{(r)}\in{\mathcal{K}}_{V(F)}(K_{s}^{(r)},{\mathcal{H}})\ \big|\ |V(K_{s}^{(r)})\cap V(F)|\geq s-r+2\big\},$
${\mathcal{K}}^{i}_{V(F)}(K_{s}^{(r)},{\mathcal{H}})=\big\{K_{s}^{(r)}\in{\mathcal{K}}_{V(F)}(K_{s}^{(r)},{\mathcal{H}})\ \big|\ |V(K_{s}^{(r)})\cap V(F)|=i\big\},$
for $1\leq i\leq s-r+1$. Clearly, $|{\mathcal{K}}^{*}_{V(F)}(K_{s}^{(r)},{\mathcal{H}})|=O(n^{r-2})$ by the definition of ${\mathcal{K}}^{*}_{V(F)}(K_{s}^{(r)},{\mathcal{H}})$. For $1\leq i\leq s-r+1$, set
${\mathcal{K}}^{i}(K_{s-i}^{(r)},{\mathcal{H}})=\big\{K_{s}^{(r)}-V(F)\ \big|\ K_{s}^{(r)}\in{\mathcal{K}}^{i}_{V(F)}(K_{s}^{(r)},{\mathcal{H}})\big\}.$
[…]
$N_{V(F)}(K_{s-i}^{(r)})=\big\{\bigcup V(K_{i}^{(r)})\ \big|\ V(K_{i}^{(r)})\subseteq V(F)\ \text{and}\ {\mathcal{H}}[V(K_{i}^{(r)})\cup V(K_{s-i}^{(r)})]=K_{s}^{(r)}\big\}.$
We now partition ${\mathcal{K}}^{i}(K_{s-i}^{(r)},{\mathcal{H}})$ into two parts for each $1\leq i\leq s-r+1$. Set
$X_{s-i}=\big\{K_{s-i}^{(r)}\in{\mathcal{K}}^{i}(K_{s-i}^{(r)},{\mathcal{H}})\ \big|\ |N_{V(F)}(K_{s-i}^{(r)})|\geq\ell^{\prime}\big\},$
$Y_{s-i}=\big\{K_{s-i}^{(r)}\in{\mathcal{K}}^{i}(K_{s-i}^{(r)},{\mathcal{H}})\ \big|\ |N_{V(F)}(K_{s-i}^{(r)})|<\ell^{\prime}\big\}.$
Then
$|{\mathcal{K}}^{i}_{V(F)}(K_{s}^{(r)},{\mathcal{H}})|\leq|X_{s-i}|\cdot\binom{|V(F)|}{i}+|Y_{s-i}|\cdot\binom{\ell^{\prime}-1}{i}$
for each $1\leq i\leq s-r+1$. It is easy to see that $|Y_{r-1}|\leq\binom{n-|V(F)|}{r-1}$ and $|Y_{s-i}|\leq{\mathcal{N}}(K_{s-i}^{(r)},{\mathcal{H}}-V(F))$ for each i<sr+1i<s-r+1. Since ${\mathcal{H}}-V(F)$ is $Q_{2}\cup\cdots\cup Q_{k}$-free, by the induction hypothesis, we have
${\mathcal{N}}\big(K_{s-i}^{(r)},{\mathcal{H}}-V(F)\big)\leq\binom{t-\ell^{\prime}}{s-i-r+1}\frac{n^{r-1}}{(r-1)!}+o(n^{r-1})$
for each $1\leq i<s-r+1$. Therefore,
$|{\mathcal{K}}^{s-r+1}_{V(F)}(K_{s}^{(r)},{\mathcal{H}})|\leq|X_{r-1}|\cdot\binom{|V(F)|}{s-r+1}+\binom{\ell^{\prime}-1}{s-r+1}\cdot\binom{n-|V(F)|}{r-1}$
and
$|{\mathcal{K}}^{i}_{V(F)}(K_{s}^{(r)},{\mathcal{H}})|\leq|X_{s-i}|\cdot\binom{|V(F)|}{i}+\binom{\ell^{\prime}-1}{i}\cdot\binom{t-\ell^{\prime}}{s-i-r+1}\frac{n^{r-1}}{(r-1)!}+o(n^{r-1})$
for each $1\leq i<s-r+1$. Thus,
$|{\mathcal{K}}_{V(F)}(K_{s}^{(r)},{\mathcal{H}})|$ $=$ $\sum_{i=1}^{s-r+1}|{\mathcal{K}}^{i}_{V(F)}(K_{s}^{(r)},{\mathcal{H}})|+|{\mathcal{K}}^{*}_{V(F)}(K_{s}^{(r)},{\mathcal{H}})|$ (5.2)
$\leq$ $\left[\sum_{i=1}^{s-r+1}|X_{s-i}|\cdot\binom{|V(F)|}{i}\right]+\Bigg[\binom{\ell^{\prime}-1}{s-r+1}\cdot\binom{n-|V(F)|}{r-1}$
$+\sum_{i=1}^{s-r}\binom{\ell^{\prime}-1}{i}\cdot\binom{t-\ell^{\prime}}{s-i-r+1}\frac{n^{r-1}}{(r-1)!}\Bigg]+o(n^{r-1})$
$\leq$ $\max_{1\leq i\leq s-r+1}\binom{|V(F)|}{i}\cdot\left(|X_{s-1}|+|X_{s-2}|+\cdots+|X_{r-1}|\right)$
$+\left(\binom{\ell^{\prime}-1}{s-r+1}+\sum_{i=1}^{s-r}\binom{\ell^{\prime}-1}{i}\cdot\binom{t-\ell^{\prime}}{s-i-r+1}\right)\frac{n^{r-1}}{(r-1)!}+o(n^{r-1})$
$=$ $\max_{1\leq i\leq s-r+1}\binom{|V(F)|}{i}\cdot\left(|X_{r-1}|+|X_{r}|+\cdots+|X_{s-1}|\right)$
$+\left(\binom{t-1}{s-r+1}-\binom{t-\ell^{\prime}}{s-r+1}\right)\frac{n^{r-1}}{(r-1)!}+o(n^{r-1}),$
where the last equality holds by Vandermonde’s identity.
Combining (5.1) and (5.2), we get
$\left(|X_{r-1}|+|X_{r}|+\cdots+|X_{s-1}|\right)>\frac{\binom{t-1}{s-r}}{\max_{1\leq i\leq s-r+1}\binom{|V(F)|}{i}}\cdot\frac{n^{r-1}}{(r-1)!}+o(n^{r-1}).$ (5.3)
By the definition of $X_{i}$ ($r-1\leq i\leq s-1$), we have $|N_{V(F)}(K_{i}^{(r)})|\geq\ell^{\prime}$ for any $K_{i}^{(r)}\in X_{i}$. Therefore, by (5.3) and the pigeonhole principle, there exists an $\ell^{\prime}$-set $U\subseteq V(F)$ such that, there are at least
$\frac{1}{\binom{|V(F)|}{\ell^{\prime}}}\cdot\frac{\binom{t-1}{s-r}}{\max_{1\leq i\leq s-r+1}\binom{|V(F)|}{i}}\cdot\frac{n^{r-1}}{(r-1)!}+o(n^{r-1})=\Omega(n^{r-1})$
copies of $K_{r-1}^{(r)},K_{r}^{(r)},\dots,K_{s-1}^{(r)}\in\bigcup_{i=r-1}^{s-1}X_{i}$ such that each $K_{i}^{(r)}$ forms a $K_{i+1}^{(r)}$ together with any vertex in $U$ (note that it is possible that $K_{i}^{(r)}\subseteq K_{j}^{(r)}$ for $r-1\leq i<j\leq s-1$). Without loss of generality, let $X^{\prime}_{i}\subseteq X_{i}$ denote the collection of such $K_{i}^{(r)}$ for each $r-1\leq i\leq s-1$, respectively. Then $\sum_{i=r-1}^{s-1}|X^{\prime}_{i}|=\Omega(n^{r-1})$. By the pigeonhole principle, $|X^{\prime}_{i}|=\Omega(n^{r-1})$ for some $r-1\leq i\leq s-1$. Note that $i$ is fixed in what follows.
Claim 5.2.
${\mathcal{H}}-U$ is $Q_{2}\cup\cdots\cup Q_{k}$-free.
Proof of Claim.
Suppose to the contrary that ${\mathcal{H}}-U$ contains a copy of $Q_{2}\cup\cdots\cup Q_{k}$, denoted by $F^{\prime}$. Therefore, $|V(F^{\prime})|=\sum_{i=2}^{k}\left(\ell_{i}(r-1)+1\right)$ if $Q_{2}\cup\cdots\cup Q_{k}=P_{\ell_{2}}^{r}\cup\cdots\cup P_{\ell_{k}}^{r}$, and $|V(F^{\prime})|=\sum_{i=2}^{k}\ell_{i}(r-1)$ if $Q_{2}\cup\cdots\cup Q_{k}=C_{\ell_{2}}^{r}\cup\cdots\cup C_{\ell_{k}}^{r}$. Moreover, $|V(F^{\prime})\backslash V(F)|\leq\sum_{i=2}^{k}\left(\ell_{i}(r-1)+1\right)$.
Observe that
$\big|\big\{K_{i}^{(r)}\in X^{\prime}_{i}\,\big|\,|V(K_{i}^{(r)})\cap V(F^{\prime})|\geq i-r+2\big\}\big|=O(n^{r-2}).$
Then there are at least $|X^{\prime}_{i}|-O(n^{r-2})=\Omega(n^{r-1})$ copies of $K_{i}^{(r)}$ in $X^{\prime}_{i}$ such that $|V(K_{i}^{(r)})\backslash V(F^{\prime})|\geq r-1$. For $r-1\leq j\leq i$, set
$X^{\prime}_{i,j}=\big\{K_{i}^{(r)}\in X^{\prime}_{i}\,\big|\,|V(K_{i}^{(r)})\backslash V(F^{\prime})|=j\big\}.$
Then $\sum_{j=r-1}^{i}|X^{\prime}_{i,j}|=\Omega(n^{r-1})$. Now we define the following collections. Let
$X^{\prime\prime}_{j}=\big\{K_{i}^{(r)}-V(F^{\prime})\,\big|\,K_{i}^{(r)}\in X^{\prime}_{i,j}\big\}$
for $r-1\leq j\leq i$. Note that $X^{\prime\prime}_{j}$ is a collection of (r1)(r-1)-sets in $V({\mathcal{H}})\backslash(V(F)\cup V(F^{\prime}))$ when j=r1j=r-1, and $X^{\prime\prime}_{j}$ is a collection of $K_{j}^{(r)}$ in $V({\mathcal{H}})\backslash(V(F)\cup V(F^{\prime}))$ when $r-1<j\leq i$. By the definitions of $X^{\prime}_{i,j}$ and $X^{\prime\prime}_{j}$, we have $|X^{\prime\prime}_{j}|\geq|X^{\prime}_{i,j}|/\binom{|V(F^{\prime})|}{i-j}$ for each $r-1\leq j\leq i$. Thus,
$\sum_{j=r-1}^{i}|X^{\prime\prime}_{j}|$ $\geq$ $\sum_{j=r-1}^{i}|X^{\prime}_{i,j}|\cdot\binom{|V(F^{\prime})|}{i-j}^{-1}$
$\geq$ $\min_{r-1\leq j\leq i}\binom{|V(F^{\prime})|}{i-j}^{-1}\cdot\left(|X^{\prime}_{i,r-1}|+|X^{\prime}_{i,r}|+\cdots+|X^{\prime}_{i,i}|\right)$
$=$ $\Omega(n^{r-1}).$
By the pigeonhole principle, $|X^{\prime\prime}_{j}|=\Omega(n^{r-1})$ for some $r-1\leq j\leq i$. Note that $j$ is also fixed in what follows.
Let ${\mathcal{S}}^{r-1}=\big\{K_{r-1}^{(r)}\ \big|\ K_{r-1}^{(r)}\subseteq V(K_{m}^{(r)})\ \text{and}\ K_{m}^{(r)}\in X^{\prime\prime}_{m},\,r-1\leq m\leq i\big\}$. Consider the (r1)(r-1)-graph ${\mathcal{H}}^{r-1}=\left(V({\mathcal{H}})\backslash(V(F)\cup V(F^{\prime})),{\mathcal{S}}^{r-1}\right)$. We claim that ${\mathcal{H}}^{r-1}$ is $P_{3\ell^{\prime}+2}^{r-1}$-free. Otherwise, assume that ${\mathcal{H}}^{r-1}$ contains a copy of $P_{3\ell^{\prime}+2}^{r-1}$. We will find a copy of $Q_{1}$ in $\mathcal{H}$} that is vertex-disjoint from $F^{\prime}$. Thus, $Q_{1}\cup F^{\prime}$ forms a copy of $Q_{1}\cup\cdots\cup Q_{k}$ in $\mathcal{H}$}, and we are done. Let $U=\{v_{1},\dots,v_{\ell^{\prime}}\}$. Recall that each $K_{i}^{(r)}\in X^{\prime}_{i}$ forms a $K_{i+1}^{(r)}$ together with any vertex in $U$. This implies that for $r-1\leq m\leq i$, each $K_{m}^{(r)}\in X^{\prime\prime}_{m}$ forms a $K_{m+1}^{(r)}$ together with any vertex in $U$. We consider the following two cases.
Case 1. Either $\ell_{1}$ is even, or $Q_{1}=P^{r}_{\ell_{1}}$.
Since $(\ell^{\prime}+1)S_{2}^{r-1}\subseteq P_{3\ell^{\prime}+2}^{r-1}$, ${\mathcal{H}}^{r-1}$ contains a copy of $(\ell^{\prime}+1)S_{2}^{r-1}$. We denote it and its hyperedges by $S^{(1)}\cup\cdots\cup S^{(\ell^{\prime}+1)}$ and $e^{(1)}_{1},e^{(1)}_{2},\dots,e^{(\ell^{\prime}+1)}_{1},e^{(\ell^{\prime}+1)}_{2}$, where $e^{(i)}_{1},e^{(i)}_{2}\in E(S^{(i)})$. Then we may construct a copy of $P^{r}_{2\ell^{\prime}}$ or a copy of $C^{r}_{2\ell^{\prime}}$ that is vertex-disjoint from $F^{\prime}$ as follows:
$P^{r}_{2\ell^{\prime}}:\ \{v_{1}\}\cup e_{2}^{(1)},\{v_{1}\}\cup e_{1}^{(2)},\{v_{2}\}\cup e_{2}^{(2)},\dots,\{v_{\ell^{\prime}}\}\cup e_{2}^{({\ell^{\prime}})},\{v_{\ell^{\prime}}\}\cup e_{1}^{({\ell^{\prime}}+1)},$
$C^{r}_{2\ell^{\prime}}:\ \{v_{1}\}\cup e_{1}^{(1)},\{v_{2}\}\cup e_{2}^{(1)},\{v_{2}\}\cup e_{1}^{(2)},\dots,\{v_{\ell^{\prime}}\}\cup e_{1}^{({\ell^{\prime}})},\{v_{1}\}\cup e_{2}^{({\ell^{\prime}})}.$
Observe that $P^{r}_{\ell_{1}}\subseteq P^{r}_{2\ell^{\prime}}$, and $C^{r}_{2\ell^{\prime}}=C^{r}_{\ell_{1}}$ when $\ell_{1}$ is even. Thus, we may find a copy of $Q_{1}$ in $\mathcal{H}$} that is vertex-disjoint from $F^{\prime}$, and we are done.
Case 2. $\ell_{1}$ is odd and $Q_{1}=C^{r}_{\ell_{1}}$.
Since $(\ell^{\prime}-2)S_{2}^{r-1}\cup P_{3}^{r-1}\subseteq P_{3\ell^{\prime}+2}^{r-1}$, ${\mathcal{H}}^{r-1}$ contains a copy of $(\ell^{\prime}-2)S_{2}^{r-1}\cup P_{3}^{r-1}$. We denote it and its hyperedges by $S^{(1)}\cup\cdots\cup S^{(\ell^{\prime}-2)}\cup P$ and $e^{(1)}_{1},e^{(1)}_{2},\dots,e^{(\ell^{\prime}-2)}_{1},e^{(\ell^{\prime}-2)}_{2},e_{1},e_{2},e_{3}$, where $e^{(i)}_{1},e^{(i)}_{2}\in E(S^{(i)})$ and $e_{1},e_{2},e_{3}\in E(P)$. Then we may construct a copy of $C^{r}_{\ell_{1}}$ that is vertex-disjoint from $F^{\prime}$ as follows:
$C^{r}_{\ell_{1}}:\ \{v_{1}\}\cup e_{1}^{(1)},\{v_{2}\}\cup e_{2}^{(1)},\dots,\{v_{\ell^{\prime}-1}\}\cup e_{2}^{({\ell^{\prime}-2})},\{v_{\ell^{\prime}-1}\}\cup e_{1},\{v_{\ell^{\prime}}\}\cup e_{2},\{v_{1}\}\cup e_{3}.$
Thus, ${\mathcal{H}}^{r-1}$ is $P_{3\ell^{\prime}+2}^{r-1}$-free. By Theorems 1.7 and 1.10, we have ${\mathcal{N}}(K_{j}^{(r-1)},{\mathcal{H}}^{r-1})=O(n^{r-2})$. By the definitions of ${\mathcal{S}}^{r-1}$ and ${\mathcal{H}}^{r-1}$, it follows that $|X^{\prime\prime}_{j}|\leq{\mathcal{N}}(K_{j}^{(r-1)},{\mathcal{H}}^{r-1})$, since each $K_{j}^{(r)}\in X^{\prime\prime}_{j}$ corresponds to a copy of $K_{j}^{(r-1)}$ in ${\mathcal{H}}^{r-1}$. Thus, $|X^{\prime\prime}_{j}|=O(n^{r-2})$, which contradicts $|X^{\prime\prime}_{j}|=\Omega(n^{r-1})$. This completes the proof of Claim 5.2. ∎
Let ${\mathcal{K}}_{U}(K_{s}^{(r)},{\mathcal{H}})=\{K_{s}^{(r)}\subseteq{\mathcal{H}}\,|\,V(K_{s}^{(r)})\cap U\neq\emptyset\}$, namely, the collection of all $K_{s}^{(r)}$ in $\mathcal{H}$} incident to vertices in $U$. Now we estimate the upper bound for $|{\mathcal{K}}_{U}(K_{s}^{(r)},{\mathcal{H}})|$. We partition ${\mathcal{K}}_{U}(K_{s}^{(r)},{\mathcal{H}})$ into the following sr+2s-r+2 parts. Let
${\mathcal{K}}^{*}_{U}(K_{s}^{(r)},{\mathcal{H}})=\big\{K_{s}^{(r)}\in{\mathcal{K}}_{U}(K_{s}^{(r)},{\mathcal{H}})\ \big|\ |V(K_{s}^{(r)})\cap U|\geq s-r+2\big\},$
${\mathcal{K}}^{i}_{U}(K_{s}^{(r)},{\mathcal{H}})=\big\{K_{s}^{(r)}\in{\mathcal{K}}_{U}(K_{s}^{(r)},{\mathcal{H}})\ \big|\ |V(K_{s}^{(r)})\cap U|=i\big\},$
for $1\leq i\leq s-r+1$. Clearly, $|{\mathcal{K}}^{*}_{U}(K_{s}^{(r)},{\mathcal{H}})|=O(n^{r-2})$ by the definition of ${\mathcal{K}}^{*}_{U}(K_{s}^{(r)},{\mathcal{H}})$.
Observe that
${\mathcal{K}}^{s-r+1}_{U}(K_{s}^{(r)},{\mathcal{H}})\leq\binom{|U|}{s-r+1}\cdot\binom{n-|U|}{r-1}=\binom{\ell^{\prime}}{s-r+1}\cdot\binom{n-\ell^{\prime}}{r-1}$
and
${\mathcal{K}}^{i}_{U}(K_{s}^{(r)},{\mathcal{H}})\leq\binom{|U|}{i}\cdot{\mathcal{N}}(K_{s-i}^{(r)},{\mathcal{H}}-U)=\binom{\ell^{\prime}}{i}\cdot{\mathcal{N}}(K_{s-i}^{(r)},{\mathcal{H}}-U)$
for $1\leq i<s-r+1$.
By Claim 5.2 and the induction hypothesis, we have
${\mathcal{N}}(K_{s-i}^{(r)},{\mathcal{H}}-U)\leq\binom{t-\ell^{\prime}}{s-i-r+1}\cdot\frac{n^{r-1}}{(r-1)!}+o(n^{r-1})$
for $1\leq i\leq s-r$. Therefore,
$|{\mathcal{K}}_{U}(K_{s}^{(r)},{\mathcal{H}})|$ $=$ $\sum_{i=1}^{s-r+1}|{\mathcal{K}}^{i}_{U}(K_{s}^{(r)},{\mathcal{H}})|+|{\mathcal{K}}^{*}_{U}(K_{s}^{(r)},{\mathcal{H}})|$
$\leq$ $\sum_{i=1}^{s-r}\binom{\ell^{\prime}}{i}\cdot{\mathcal{N}}(K_{s-i}^{(r)},{\mathcal{H}}-U)+\binom{\ell^{\prime}}{s-r+1}\cdot\binom{n-\ell^{\prime}}{r-1}+O(n^{r-2})$
$\leq$ $\sum_{i=1}^{s-r}\binom{\ell^{\prime}}{i}\cdot\left(\binom{t-\ell^{\prime}}{s-i-r+1}\cdot\frac{n^{r-1}}{(r-1)!}+o(n^{r-1})\right)+\binom{\ell^{\prime}}{s-r+1}\cdot\binom{n-\ell^{\prime}}{r-1}+O(n^{r-2})$
$=$ $\left(\sum_{i=1}^{s-r}\binom{\ell^{\prime}}{i}\cdot\binom{t-\ell^{\prime}}{s-i-r+1}+\binom{\ell^{\prime}}{s-r+1}\right)\frac{n^{r-1}}{(r-1)!}+o(n^{r-1})$
$=$ $\left(\binom{t}{s-r+1}-\binom{t-\ell^{\prime}}{s-r+1}\right)\frac{n^{r-1}}{(r-1)!}+o(n^{r-1}),$
where the last equality holds by Vandermonde’s identity.
Thus,
${\mathcal{N}}(K_{s}^{(r)},{\mathcal{H}}-U)$ $=$ ${\mathcal{N}}(K_{s}^{(r)},{\mathcal{H}})-|{\mathcal{K}}_{U}(K_{s}^{(r)},{\mathcal{H}})|$
$>$ $\binom{t}{s-r+1}\frac{n^{r-1}}{(r-1)!}-\left(\binom{t}{s-r+1}-\binom{t-\ell^{\prime}}{s-r+1}\right)\frac{n^{r-1}}{(r-1)!}+o(n^{r-1})$
$=$ $\binom{t-\ell^{\prime}}{s-r+1}\frac{n^{r-1}}{(r-1)!}+o(n^{r-1}).$
This implies that ${\mathcal{H}}-U$ contains a copy of $Q_{2}\cup\cdots\cup Q_{k}$ by the induction hypothesis, which contradicts Claim 5.2. This completes the proof. ∎
We now begin with the proof of Theorem 2.10 that we restate here for convenience.
Theorem 2.10.
Let integers $r\geq 3$, $k\geq 1$, $p\geq 0$, $\ell_{1},\dots,\ell_{k}\geq 5$, $t_{1},\dots,t_{p}\geq 2$ if $p\geq 1$, and $t=\sum_{i=1}^{k}\lfloor\frac{\ell_{i}+1}{2}\rfloor+p-1$. Suppose $Q_{i}\in\{P_{\ell_{i}}^{r},C_{\ell_{i}}^{r}\}$, $Q_{1}\cup\cdots\cup Q_{k}\in\{P_{\ell_{1}}^{r}\cup\cdots\cup P_{\ell_{k}}^{r},C_{\ell_{1}}^{r}\cup\cdots\cup C_{\ell_{k}}^{r}\}$ and $n$ is sufficiently large.
(i) If $s\leq t+r-1$, then
${\rm{ex}}_{r}(n,K_{s}^{(r)},Q_{1}\cup\cdots\cup Q_{k}\cup S^{r}_{t_{1}}\cup\cdots\cup S^{r}_{t_{p}})=(1+o(1)){\mathcal{N}}(K_{s}^{(r)},S_{n,t}^{r})=\binom{t}{s-r+1}\frac{n^{r-1}}{(r-1)!}+o(n^{r-1}).$
(ii) If $s\geq t+r$, then
${\rm{ex}}_{r}(n,K_{s}^{(r)},Q_{1}\cup\cdots\cup Q_{k}\cup S^{r}_{t_{1}}\cup\cdots\cup S^{r}_{t_{p}})=o(n^{r-1}).$
Proof..
For the lower bound, we consider the $r$-graph $S_{n,t}^{r}$. Observe that any $P_{\ell_{i}}^{r}$ or $C_{\ell_{i}}^{r}$ in $S_{n,t}^{r}$ contains at least $\big\lfloor\frac{\ell_{i}+1}{2}\big\rfloor$ vertices in the $t$-set of $S_{n,t}^{r}$, and any $S_{t_{i}}^{r}$ in $S_{n,t}^{r}$ contains at least one vertex in the $t$-set of $S_{n,t}^{r}$. Thus, $S_{n,t}^{r}$ is $Q_{1}\cup\cdots\cup Q_{k}\cup S^{r}_{t_{1}}\cup\cdots\cup S^{r}_{t_{p}}$-free as $t<\sum_{i=1}^{k}\lfloor\frac{\ell_{i}+1}{2}\rfloor+p$.
We now prove the upper bounds in (i) and (ii) together. Suppose $\mathcal{H}$} is an $r$-graph on $n$ vertices with
${\mathcal{N}}(K_{s}^{(r)},{\mathcal{H}})>\binom{t}{s-r+1}\frac{n^{r-1}}{(r-1)!}+o(n^{r-1}).$
This implies that ${\mathcal{N}}(K_{s}^{(r)},{\mathcal{H}})\sim an^{r-1}$ as $n$ is sufficiently large, where $a>\binom{t}{s-r+1}\frac{1}{(r-1)!}$ if $s\leq t+r-1$, and $a>0$ if $s\geq t+r$. Next, we will show that $\mathcal{H}$} contains a copy of $Q_{1}\cup\cdots\cup Q_{k}\cup S^{r}_{t_{1}}\cup\cdots\cup S^{r}_{t_{p}}$ by induction on $p$. If $p=0$, then the conclusion follows from Theorem 2.8. Suppose $p\geq 1$ and the statement holds for any $p^{\prime}<p$. By Theorem 2.5, there exists a constant $c(s,r,t_{p})$ such that
${\rm{ex}}_{r}(n,K_{s}^{(r)},S_{t_{p}}^{r})\leq c(s,r,t_{p})n^{r-2}.$ (5.4)
Let $u\in V({\mathcal{H}})$ be a vertex such that $|{\mathcal{K}}_{u}(K_{s}^{(r)},{\mathcal{H}})|=\max_{v\in V({\mathcal{H}})}|{\mathcal{K}}_{v}(K_{s}^{(r)},{\mathcal{H}})|$. Assume that $|{\mathcal{K}}_{u}(K_{s}^{(r)},{\mathcal{H}})|<\frac{{\mathcal{N}}(K_{s}^{(r)},{\mathcal{H}})-c(s,r,t_{p})n^{r-2}}{\sum_{i=1}^{k}\left(\ell_{i}(r-1)+1\right)+\sum_{i=1}^{p-1}\left(t_{i}(r-1)+1\right)}$. Since $n$ is sufficiently large, we have ${\mathcal{N}}(K_{s}^{(r)},{\mathcal{H}})\sim an^{r-1}>\binom{t-1}{s-r+1}\frac{n^{r-1}}{(r-1)!}+o(n^{r-1})$. By induction hypothesis, $\mathcal{H}$} contains a copy of $Q_{1}\cup\cdots\cup Q_{k}\cup S^{r}_{t_{1}}\cup\cdots\cup S^{r}_{t_{p-1}}$, denoted by $Q$. Therefore, $|V(Q)|=\sum_{i=1}^{k}\left(\ell_{i}(r-1)+1\right)+\sum_{i=1}^{p-1}\left(t_{i}(r-1)+1\right)$. Then
${\mathcal{N}}(K_{s}^{(r)},{\mathcal{H}}-V(Q))$ $>$ ${\mathcal{N}}(K_{s}^{(r)},{\mathcal{H}})-|V(Q)|\cdot\frac{{\mathcal{N}}(K_{s}^{(r)},{\mathcal{H}})-c(s,r,t_{p})n^{r-2}}{\sum_{i=1}^{k}\left(\ell_{i}(r-1)+1\right)+\sum_{i=1}^{p-1}\left(t_{i}(r-1)+1\right)}$
$=$ $c(s,r,t_{p})n^{r-2},$
which implies ${\mathcal{H}}-V(Q)$ contains a copy of $S_{t_{p}}^{r}$ by (5.4). Thus, $\mathcal{H}$} contains a copy of $Q_{1}\cup\cdots\cup Q_{k}\cup S^{r}_{t_{1}}\cup\cdots\cup S^{r}_{t_{p}}$, and we are done.
Now suppose $|{\mathcal{K}}_{u}(K_{s}^{(r)},{\mathcal{H}})|\geq\frac{{\mathcal{N}}(K_{s}^{(r)},{\mathcal{H}})-c(s,r,t_{p})n^{r-2}}{\sum_{i=1}^{k}\left(\ell_{i}(r-1)+1\right)+\sum_{i=1}^{p-1}\left(t_{i}(r-1)+1\right)}$. Let
${\mathcal{K}}^{s-1}=\big\{K_{s}^{(r)}-\{u\}\ \big|\ K_{s}^{(r)}\in{\mathcal{K}}_{u}(K_{s}^{(r)},{\mathcal{H}})\big\}.$
Then $|{\mathcal{K}}^{s-1}|=|{\mathcal{K}}_{u}(K_{s}^{(r)},{\mathcal{H}})|=\Omega(n^{r-1})$. Similar to the proof of Claim 4.2, we may claim that ${\mathcal{H}}-\{u\}$ is $Q_{1}\cup\cdots\cup Q_{k}\cup S^{r}_{t_{1}}\cup\cdots\cup S^{r}_{t_{p-1}}$-free. By the induction hypothesis,
${\mathcal{N}}(K_{s-1}^{(r)},{\mathcal{H}}-\{u\})\leq\binom{t-1}{s-r}\frac{n^{r-1}}{(r-1)!}+o(n^{r-1}).$
Therefore,
$|{\mathcal{K}}_{u}(K_{s}^{(r)},{\mathcal{H}})|\leq{\mathcal{N}}(K_{s-1}^{(r)},{\mathcal{H}}-\{u\})\leq\binom{t-1}{s-r}\frac{n^{r-1}}{(r-1)!}+o(n^{r-1}).$
Thus,
${\mathcal{N}}(K_{s}^{(r)},{\mathcal{H}}-\{u\})$ $=$ ${\mathcal{N}}(K_{s}^{(r)},{\mathcal{H}})-|{\mathcal{K}}_{u}(K_{s}^{(r)},{\mathcal{H}})|$
$>$ $\binom{t}{s-r+1}\frac{n^{r-1}}{(r-1)!}-\binom{t-1}{s-r}\frac{n^{r-1}}{(r-1)!}+o(n^{r-1})$
$=$ $\binom{t-1}{s-r+1}\frac{n^{r-1}}{(r-1)!}+o(n^{r-1}).$
This implies that ${\mathcal{H}}-\{u\}$ contains a copy of $Q_{1}\cup\cdots\cup Q_{k}\cup S^{r}_{t_{1}}\cup\cdots\cup S^{r}_{t_{p-1}}$ by the induction hypothesis, which is a contradiction. This completes the proof. ∎
6 Concluding remarks
6.1 Summary of main results
In this work, we initiate a systematic study of generalized Turán problems for expansions. Our study considers both non-degenerate and degenerate settings and obtains several exact and asymptotic results. More specifically, in the non-degenerate case, we show that generalized Turán problems for expansions of graphs $F$ are non-degenerate if and only if $\chi(F)>s$ (see Proposition 2.1). Moreover, we determine the exact generalized Turán numbers for expansions of complete graphs (see Theorem 2.2), and further establish the asymptotics of the generalized Turán number for expansions of the vertex-disjoint union of complete graphs (see Theorem 2.4). In the degenerate case, we determine the order of magnitude of $\mathrm{ex}_{r}(n,K_{s}^{(r)},S_{\ell}^{r})$ (see Theorem 2.5), and establish the asymptotics of generalized Turán numbers for expansions of several classes of forests, including star forests (see Theorem 2.6), linear forests (see Theorem 2.8) and star-path forests (see Theorem 2.10).
$\bullet$ In Theorem 2.8, we consider the cases where $Q_{1}\cup\cdots\cup Q_{k}\in\{P_{\ell_{1}}^{r}\cup\cdots\cup P_{\ell_{k}}^{r},C_{\ell_{1}}^{r}\cup\cdots\cup C_{\ell_{k}}^{r}\}$ with $Q_{i}\in\{P_{\ell_{i}}^{r},C_{\ell_{i}}^{r}\}$ for each $i$. By a slight modification of the proof of Theorem 2.8, we may obtain the following more general result on expansions of vertex-disjoint unions of paths and cycles.
Theorem 6.1.
Let $r\geq 3$, $k\geq 1$ and $\ell_{1},\dots,\ell_{k}\geq 5$ be integers and $t=\sum_{i=1}^{k}\lfloor\frac{\ell_{i}+1}{2}\rfloor-1$. Suppose $Q_{i}\in\{P_{\ell_{i}}^{r},C_{\ell_{i}}^{r}\}$ and $n$ is sufficiently large.
(i) If $s\leq t+r-1$, then
${\rm{ex}}_{r}(n,K_{s}^{(r)},Q_{1}\cup\cdots\cup Q_{k})=(1+o(1)){\mathcal{N}}(K_{s}^{(r)},S_{n,t}^{r})=\binom{t}{s-r+1}\frac{n^{r-1}}{(r-1)!}+o(n^{r-1}).$
(ii) If $s\geq t+r$, then
${\rm{ex}}_{r}(n,K_{s}^{(r)},Q_{1}\cup\cdots\cup Q_{k})=o(n^{r-1}).$
$\bullet$ The proof of Theorem 2.10 proceeds by induction on $p$, with the base case given by Theorem 2.8. Following the same inductive approach, and using Theorem 6.1 as the base case, we may obtain the following result concerning expansions of vertex-disjoint unions of paths, cycles and stars.
Theorem 6.2.
Let integers $r\geq 3$, $k\geq 1$, $p\geq 0$, $\ell_{1},\dots,\ell_{k}\geq 5$, $t_{1},\dots,t_{p}\geq 2$ if $p\geq 1$, and $t=\sum_{i=1}^{k}\lfloor\frac{\ell_{i}+1}{2}\rfloor+p-1$. Suppose $Q_{i}\in\{P_{\ell_{i}}^{r},C_{\ell_{i}}^{r}\}$ and $n$ is sufficiently large.
(i) If $s\leq t+r-1$, then
${\rm{ex}}_{r}(n,K_{s}^{(r)},Q_{1}\cup\cdots\cup Q_{k}\cup S^{r}_{t_{1}}\cup\cdots\cup S^{r}_{t_{p}})=(1+o(1)){\mathcal{N}}(K_{s}^{(r)},S_{n,t}^{r})=\binom{t}{s-r+1}\frac{n^{r-1}}{(r-1)!}+o(n^{r-1}).$
(ii) If $s\geq t+r$, then
${\rm{ex}}_{r}(n,K_{s}^{(r)},Q_{1}\cup\cdots\cup Q_{k}\cup S^{r}_{t_{1}}\cup\cdots\cup S^{r}_{t_{p}})=o(n^{r-1}).$
$\bullet$ In Theorems 2.8 and 2.10, we establish only the asymptotics of the corresponding generalized Turán numbers, since our arguments rely on an inductive argument whose base case is an asymptotic result of Axenovich, Gerbner, Liu, and Patk s [2] (see Theorem 1.10). Consequently, if the exact generalized Turán numbers for $P_{\ell}^{r}$ and $C_{\ell}^{r}$ were known, our method may yield the exact generalized Turán numbers for $Q_{1}\cup\cdots\cup Q_{k}$ and $Q_{1}\cup\cdots\cup Q_{k}\cup S^{r}_{t_{1}}\cup\cdots\cup S^{r}_{t_{p}}$, where $Q_{i}\in\{P_{\ell_{i}}^{r},C_{\ell_{i}}^{r}\}$.
6.2 A general bound on $\mathrm{ex}_{r}(n,K_{s}^{(r)},{\mathcal{F}})$
In Section 3, we establish a general upper bound on ${\rm{ex}}_{r}(n,K_{s}^{(r)},F^{r})$ for an arbitrary graph $F$ (see Lemma 3.1). Theorems 1.1 and 2.2 show that this bound is sharp in a specific case. Next, we present a more general bound on $\mathrm{ex}_{r}(n,K_{s}^{(r)},{\mathcal{F}})$ for any $r$-graph $\mathcal{F}$}. Given a graph $F$, an $r$-graph $\mathcal{H}$} is a Berge-$F$ if there is a bijection $\phi:E(F)\rightarrow E({\mathcal{H}})$ such that $e\subseteq\phi(e)$ for each $e\in E(F)$. Berge [4] defined the Berge-cycle, and Győri, Katona and Lemons [34] defined the Berge-path. Later, Gerbner and Palmer [31] extended the definition to arbitrary graphs $F$. A brief survey of Turán-type results for Berge-graphs can be found in Subsection 5.2.2 in [33]. Here, we further generalize the definition of Berge-graphs to arbitrary $r$-graphs. For $s\geq r$ and an $r$-graph $\mathcal{F}$}, an $s$-graph $\mathcal{G}$} is a Berge-$\mathcal{F}$} if there is a bijection $\phi:E({\mathcal{F}})\rightarrow E({\mathcal{G}})$ such that $e\subseteq\phi(e)$ for each $e\in E({\mathcal{F}})$. Note that for a fixed $r$-graph $\mathcal{F}$}, there are many $s$-graphs that form a Berge-$\mathcal{F}$}. For convenience, we refer to this collection of $s$-graphs as “Berge-$\mathcal{F}$}”. The maximum number of hyperedges in an $n$-vertex $s$-graph containing no subhypergraph isomorphic to any Berge-$\mathcal{F}$} is denoted by ${\rm{ex}}_{s}(n,{\rm Berge}{\text{-}}{\mathcal{F}})$.
The problem of counting copies of the complete $r$-graph $K_{s}^{(r)}$ in an $n$-vertex $\mathcal{F}$}-free $r$-graph is closely related to the study of Berge-hypergraphs. In what follows, we establish an explicit relationship between these two classes of problems.
Proposition 6.3.
Let $s\geq r\geq 2$ be integers and $\mathcal{F}$} be an $r$-graph. Then
${\rm{ex}}_{s}(n,{\rm Berge}{\text{-}}{\mathcal{F}})-\mathrm{ex}_{r}(n,{\mathcal{F}})\leq\mathrm{ex}_{r}(n,K_{s}^{(r)},{\mathcal{F}})\leq{\rm{ex}}_{s}(n,{\rm Berge}{\text{-}}{\mathcal{F}}).$
Proof..
Let $\mathcal{H}$} be an $n$-vertex $\mathcal{F}$}-free $r$-graph. It is easy to see that the $s$-clique hypergraph of $\mathcal{H}$} must not contain a Berge-$\mathcal{F}$}. This implies that $\mathrm{ex}_{r}(n,K_{s}^{(r)},{\mathcal{F}})\leq{\rm{ex}}_{s}(n,{\rm Berge}{\text{-}}{\mathcal{F}})$.
Now we consider the lower bound. Let $\mathcal{G}$} be an $n$-vertex $s$-graph containing no subhypergraph isomorphic to any Berge-$\mathcal{F}$}, and let $E({\mathcal{G}})=\{e_{1},e_{2},\dots,e_{x}\}$ denote its hyperedge set. We construct an $r$-graph $\mathcal{H}$} on the vertex set $V({\mathcal{G}})$ as follows. At step $i$ ($1\leq i\leq x$), we attempt to choose an $r$-subset of $e_{i}$ that has not yet been selected as a hyperedge of $\mathcal{H}$}. If no such $r$-subset exists, that is, if every $r$-subset of $e_{i}$ is already a hyperedge of $\mathcal{H}$}, then no hyperedge is added at this step.
Any copy of $\mathcal{F}$} in $\mathcal{H}$} would correspond to a Berge-$\mathcal{F}$} in $\mathcal{G}$}. Therefore, $\mathcal{H}$} is $\mathcal{F}$}-free. By the definition of $\mathcal{H}$}, for each hyperedge $e_{i}\in E({\mathcal{G}})$ for which no $r$-subset was added to $\mathcal{H}$}, the vertices of $e_{i}$ span a copy of $K_{s}^{(r)}$ in $\mathcal{H}$}. Consequently, the number of such hyperedges of $\mathcal{G}$} is at most the number of copies of $K_{s}^{(r)}$ in $\mathcal{H}$}. Hence, each hyperedge $e_{i}\in E({\mathcal{G}})$ corresponds either to a hyperedge of $\mathcal{H}$} or to a copy of $K_{s}^{(r)}$ spanned by its vertices in $\mathcal{H}$}. This implies that $e({\mathcal{G}})\leq\mathrm{ex}_{r}(n,K_{s}^{(r)},{\mathcal{F}})+\mathrm{ex}_{r}(n,{\mathcal{F}})$. Thus, ${\rm{ex}}_{s}(n,{\rm Berge}{\text{-}}{\mathcal{F}})-\mathrm{ex}_{r}(n,{\mathcal{F}})\leq\mathrm{ex}_{r}(n,K_{s}^{(r)},{\mathcal{F}})$. ∎
Proposition 6.3 immediately yields the following corollary, which was previously established by Gerbner and Palmer [32].
Corollary 6.4 (Gerbner and Palmer [32]).
Let $s\geq 2$ be an integer and $F$ be a graph. Then
$\mathrm{ex}(n,K_{s},F)\leq{\rm{ex}}_{s}(n,{\rm Berge}{\text{-}}F)\leq\mathrm{ex}(n,K_{s},F)+\mathrm{ex}(n,F).$
Funding
The research of Zhou and Yuan was supported by the National Natural Science Foundation of China (Nos. 12271337 and 12371347).
The research of Zhao was supported by the China Scholarship Council (No. 202506210250).
Declaration of interest
The authors declare no known conflicts of interest.
Appendix A Proof of Theorem 5.1
First, we state the following summary result regarding expansions of paths and cycles, which will be used in the proof.
Theorem A.1 (Füredi, Jiang, Seiver [26], Kostochka, Mubayi, Verstraëte [38], Frankl and Füredi [23]).
Let $r\geq 3$ and $\ell\geq 3$. For sufficiently large $n$,
${\rm{ex}}_{r}(n,P_{\ell}^{r})=\binom{n}{r}-\binom{n-\lfloor\frac{\ell-1}{2}\rfloor}{r}+\begin{cases}0&{\rm{if}}\ \ell\ {\rm{is}}\ {\rm{odd}},\\ \binom{n-\lfloor\frac{\ell-1}{2}\rfloor-2}{r-2}&{\rm{if}}\ \ell\ {\rm{is}}\ {\rm{even}}.\end{cases}$
The same result holds for $C_{\ell}^{r}$ when $r\geq 3$ and $\ell\geq 4$, except the case $(r,\ell)=(3,4)$, in which case
${\rm{ex}}_{3}(n,C_{4}^{3})=\binom{n}{3}-\binom{n-1}{3}+\max\left\{n-3,4\left\lfloor\frac{n-1}{4}\right\rfloor\right\}.$
Observe that ${\rm{ex}}_{r}(n,P^{r}_{3})={\rm{ex}}_{r}(n,K^{r}_{3})={\rm{ex}}_{r}(n,C^{r}_{3})$ for sufficiently large $n$. Now let us start with the proof of Theorem 5.1 that we restate here for convenience.
Theorem A.2 (Zhou and Yuan [51]).
Let integers $r\geq 3$, $k\geq 1$, $0\leq p\leq k$ and $\ell_{1},\dots,\ell_{k}\geq 3$, where $\ell_{i}\neq 4$ for all $i$ if $r=3$. Then for sufficiently large $n$,
${\rm{ex}}_{r}(n,P_{\ell_{1}}^{r}\cup\cdots\cup P_{\ell_{p}}^{r}\cup C_{\ell_{p+1}}^{r}\cup\cdots\cup C_{\ell_{k}}^{r})=\binom{n}{r}-\binom{n-\sum_{i=1}^{k}\big\lfloor\frac{\ell_{i}+1}{2}\big\rfloor+1}{r}+c(n;\ell_{1},\dots,\ell_{k}),$
where $c(n;\ell_{1},\dots,\ell_{k})=\binom{n-\sum_{i=1}^{k}\lfloor\frac{\ell_{i}+1}{2}\rfloor-1}{r-2}$ if all of $\ell_{1},\dots,\ell_{k}$ are even, and $c(n;\ell_{1},\dots,\ell_{k})=0$ otherwise.
Proof..
For the lower bound, we consider the $r$-graph on $n$ vertices consisting of all hyperedges intersecting a fixed set $U$ of $\sum_{i=1}^{k}\big\lfloor\frac{\ell_{i}+1}{2}\big\rfloor-1$ vertices, along with all hyperedges disjoint from $U$ containing two fixed vertices outside $U$ when all of $\ell_{1},\dots,\ell_{k}$ are even. Observe that any $P_{\ell_{i}}^{r}$ or $C_{\ell_{i}}^{r}$ in the resulting $r$-graph contains at least $\big\lfloor\frac{\ell_{i}+1}{2}\big\rfloor$ vertices in $U$. Thus, the resulting $r$-graph is $\mathcal{G}_{1}$-free as $|U|<\sum_{i=1}^{k}\big\lfloor\frac{\ell_{i}+1}{2}\big\rfloor$. Clearly, the number of hyperedges of the resulting $r$-graph is $\binom{n}{r}-\binom{n-\sum_{i=1}^{k}\lfloor\frac{\ell_{i}+1}{2}\rfloor+1}{r}+c(n;\ell_{1},\dots,\ell_{k})$.
We now proceed with the upper bound. Let $\mathcal{H}$} be an $r$-graph on $n$ vertices with
$e({\mathcal{H}})>\binom{n}{r}-\binom{n-\sum_{i=1}^{k}\big\lfloor\frac{\ell_{i}+1}{2}\big\rfloor+1}{r}+c(n;\ell_{1},\dots,\ell_{k}).$
We will show that $\mathcal{H}$} contains a copy of $P_{\ell_{1}}^{r}\cup\cdots\cup P_{\ell_{p}}^{r}\cup C_{\ell_{p+1}}^{r}\cup\cdots\cup C_{\ell_{k}}^{r}$ by induction on $k$. If $k=1$, then $e({\mathcal{H}})>{\rm{ex}}_{r}(n,P^{r}_{\ell_{1}})={\rm{ex}}_{r}(n,C^{r}_{\ell_{1}})$ by Theorem A.1, and hence the conclusion holds. Suppose $k\geq 2$ and the statement holds for any $k^{\prime}<k$. Since it is possible that $p=0$, for convenience we denote $P_{\ell_{1}}^{r}\cup\cdots\cup P_{\ell_{p}}^{r}\cup C_{\ell_{p+1}}^{r}\cup\cdots\cup C_{\ell_{k}}^{r}$ by $F_{\ell_{1}}^{r}\cup\cdots\cup F_{\ell_{k}}^{r}$.
Since $n$ is sufficiently large, we have
$e({\mathcal{H}})>\binom{n}{r}-\binom{n-\sum_{i=1}^{k}\big\lfloor\frac{\ell_{i}+1}{2}\big\rfloor+1}{r}>\binom{n}{r}-\binom{n-\big\lfloor\frac{\ell_{1}+1}{2}\big\rfloor+1}{r}+\binom{n-\big\lfloor\frac{\ell_{1}+1}{2}\big\rfloor-1}{r-2}$,
which implies $\mathcal{H}$} contains a copy of $F_{\ell_{1}}^{r}$ by Theorem A.1, denoted by $F$. Therefore, $\ell_{1}(r-1)\leq|V(F)|\leq\ell_{1}(r-1)+1$. We may assume that
$e({\mathcal{H}}-V(F))\leq\binom{n-|V(F)|}{r}-\binom{n-|V(F)|-\sum_{i=2}^{k}\big\lfloor\frac{\ell_{i}+1}{2}\big\rfloor+1}{r}+c(n-|V(F)|;\ell_{2},\dots,\ell_{k}).$
Otherwise, by induction hypothesis, we may find a copy of $F_{\ell_{2}}^{r}\cup\cdots\cup F_{\ell_{k}}^{r}$ in ${\mathcal{H}}-V(F)$. Thus, $H$ contains a copy of $F_{\ell_{1}}^{r}\cup\cdots\cup F_{\ell_{k}}^{r}$ and we are done.
Let $E_{V(F)}$ denote the set of hyperedges of $\mathcal{H}$} incident to vertices in $V(F)$. Then
$|E_{V(F)}|$ $=$ $e({\mathcal{H}})-e({\mathcal{H}}-V(F))$ (A.1)
$>$ $\binom{n}{r}-\binom{n-\sum_{i=1}^{k}\big\lfloor\frac{\ell_{i}+1}{2}\big\rfloor+1}{r}+c(n;\ell_{1},\dots,\ell_{k})$
$-\bigg[\binom{n-|V(F)|}{r}-\binom{n-|V(F)|-\sum_{i=2}^{k}\big\lfloor\frac{\ell_{i}+1}{2}\big\rfloor+1}{r}+c(n-|V(F)|;\ell_{2},\dots,\ell_{k})\bigg]$
$=$ $\frac{\lfloor\frac{\ell_{1}+1}{2}\big\rfloor}{(r-1)!}n^{r-1}+o(n^{r-1}).$
[…]
Note that the number of hyperedges in $E_{V(F)}$ containing at least two vertices from $V(F)$ is at most $\binom{|V(F)|}{2}\binom{n-2}{r-2}$. Then
$|E_{V(F)}|$ $\leq$ $|X|\cdot|V(F)|+|Y|\cdot\Big(\Big\lfloor\frac{\ell_{1}+1}{2}\Big\rfloor-1\Big)+\binom{|V(F)|}{2}\binom{n-2}{r-2}$ (A.2)
$\leq$ $|X|\cdot|V(F)|+\binom{n-|V(F)|}{r-1}\Big(\Big\lfloor\frac{\ell_{1}+1}{2}\Big\rfloor-1\Big)+\binom{|V(F)|}{2}\binom{n-2}{r-2}$.
Combining (A.1) and (A.2), we get
$|X|\geq\frac{n^{r-1}}{|V(F)|\cdot(r-1)!}+o(n^{r-1}).$ (A.3)
By the definition of $X$, for any $S\in X$ there is at least $\big\lfloor\frac{\ell_{1}+1}{2}\big\rfloor$ vertices in $V(F)$ such that it forms a hyperedge in $E_{V(F)}^{1}(S)$ together with $S$. Therefore, by (A.3) and pigeonhole principle, there exists a $\big\lfloor\frac{\ell_{1}+1}{2}\big\rfloor$-set $U$ in $V(F)$ such that, there are at least
$\frac{n^{r-1}}{|V(F)|\cdot(r-1)!}\cdot\frac{1}{\binom{|V(F)|}{\lfloor\frac{\ell_{1}+1}{2}\rfloor}}+o(n^{r-1})\geq\frac{n^{r-1}}{(\ell_{1}(r-1)+1)\cdot(r-1)!}\cdot\frac{1}{\binom{(\ell_{1}(r-1)+1)}{\lfloor\frac{\ell_{1}+1}{2}\rfloor}}+o(n^{r-1})$
[…]
$|\mathcal{S}_{U}|=\Omega(n^{r-1}).$ (A.4)
Let $E_{U}$ denote the set of hyperedges of $\mathcal{H}$} incident to vertices in $U$. Then
$|E_{U}|$ $\leq$ $\binom{|U|}{r}+\binom{|U|}{r-1}\binom{n-|U|}{1}+\cdots+\binom{|U|}{1}\binom{n-|U|}{r-1}$
$=$ $\binom{n}{r}-\binom{n-|U|}{r}=\binom{n}{r}-\binom{n-\big\lfloor\frac{\ell_{1}+1}{2}\big\rfloor}{r}$,
where the second equality holds by Vandermonde’s identity.
Now we compare the values of $c(n;\ell_{1},\dots,\ell_{k})$ and $c(n-|U|;\ell_{2},\dots,\ell_{k})$.
  • If all of $\ell_{1},\dots,\ell_{k}$ are even, then $c(n;\ell_{1},\dots,\ell_{k})=c(n-|U|;\ell_{2},\dots,\ell_{k})=\binom{n-\sum_{i=1}^{k}\big\lfloor\frac{\ell_{i}+1}{2}\big\rfloor-1}{r-2}$.
  • If at least two of $\ell_{1},\dots,\ell_{k}$ are odd, then $c(n;\ell_{1},\dots,\ell_{k})=c(n-|U|;\ell_{2},\dots,\ell_{k})=0$.
  • If exactly one of $\ell_{1},\dots,\ell_{k}$ is odd and $\ell_{1}$ is even, then $c(n;\ell_{1},\dots,\ell_{k})=c(n-|U|;\ell_{2},\dots,\ell_{k})=0$.
  • If exactly one of $\ell_{1},\dots,\ell_{k}$ is odd and $\ell_{1}$ is odd, then we exchange $F^{r}_{\ell_{1}}$ and $F^{r}_{\ell_{2}}$, and apply arguments similar to those above to $F^{r}_{\ell^{\prime}_{1}}\cup F^{r}_{\ell^{\prime}_{2}}\cup\cdots\cup F^{r}_{\ell_{k}}$, where $F^{r}_{\ell^{\prime}_{1}}=F^{r}_{\ell_{2}}$ and $F^{r}_{\ell^{\prime}_{2}}=F^{r}_{\ell_{1}}$. Therefore, one of $\ell^{\prime}_{1},\ell^{\prime}_{2},\dots,\ell_{k}$ is odd and $\ell^{\prime}_{1}$ is even. It follows that $c(n;\ell^{\prime}_{1},\ell^{\prime}_{2},\dots,\ell_{k})=c(n-|U|;\ell^{\prime}_{2},\dots,\ell_{k})=0$.
Then $c(n;\ell_{1},\dots,\ell_{k})-c(n-|U|;\ell_{2},\dots,\ell_{k})\geq 0$ always holds. Therefore,
$\binom{n}{r}-\binom{n-\sum_{i=1}^{k}\big\lfloor\frac{\ell_{i}+1}{2}\big\rfloor+1}{r}+c(n;\ell_{1},\dots,\ell_{k})$
$-\bigg[\binom{n-\big\lfloor\frac{\ell_{1}+1}{2}\big\rfloor}{r}-\binom{n-\big\lfloor\frac{\ell_{1}+1}{2}\big\rfloor-\sum_{i=2}^{k}\big\lfloor\frac{\ell_{i}+1}{2}\big\rfloor+1}{r}+c(n-|U|;\ell_{2},\dots,\ell_{k})\bigg]$
$=$ $\binom{n}{r}-\binom{n-\big\lfloor\frac{\ell_{1}+1}{2}\big\rfloor}{r}+c(n;\ell_{1},\dots,\ell_{k})-c(n-|U|;\ell_{2},\dots,\ell_{k})$
$\geq$ $|E_{U}|.$
Namely,
$e({\mathcal{H}}-U)$ $=$ $e({\mathcal{H}})-|E_{U}|$
$>$ $\binom{n}{r}-\binom{n-\sum_{i=1}^{k}\big\lfloor\frac{\ell_{i}+1}{2}\big\rfloor+1}{r}+c(n;\ell_{1},\dots,\ell_{k})-|E_{U}|$
$\geq$ $\binom{n-\big\lfloor\frac{\ell_{1}+1}{2}\big\rfloor}{r}-\binom{n-\big\lfloor\frac{\ell_{1}+1}{2}\big\rfloor-\sum_{i=2}^{k}\big\lfloor\frac{\ell_{i}+1}{2}\big\rfloor+1}{r}+c(n-|U|;\ell_{2},\dots,\ell_{k}),$
[…]
Next, we will find a copy of $F^{r}_{\ell_{1}}$ that is vertex-disjoint from $V(F^{\prime})$ in $\mathcal{H}$}. Therefore, $F^{r}_{\ell_{1}}$ and $F^{\prime}$ together form a copy of $F^{r}_{\ell_{1}}\cup\cdots\cup F^{r}_{\ell_{k}}$ in $\mathcal{H}$}. Consider the (r1)(r-1)-graph ${\mathcal{H}}^{r-1}=(V({\mathcal{H}})\backslash U,\mathcal{S}_{U})$. Note that the number of hyperedges in $E({\mathcal{H}}^{r-1})$ that contain at least one vertex from $V(F^{\prime})$ is at most $|V(F^{\prime})|\cdot\binom{n-|U|-1}{r-2}\leq\sum_{i=2}^{k}[(r-1)\ell_{i}+1]\cdot\binom{n-|U|-1}{r-2}$. By (A.4) and Theorem 1.3, we have
$e({\mathcal{H}}^{r-1}[V({\mathcal{H}})\backslash(U\cup V(F^{\prime}))])$ $\geq$ $e({\mathcal{H}}^{r-1})-\sum_{i=2}^{k}\big[(r-1)\ell_{i}+1\big]\cdot\binom{n-|U|-1}{r-2}$
$=$ $|\mathcal{S}_{U}|-\sum_{i=2}^{k}\big[(r-1)\ell_{i}+1\big]\cdot\binom{n-|U|-1}{r-2}$
$=$ $\Omega(n^{r-1})$
$>$ ${\rm{ex}}_{r-1}(n-|U|,S_{2}^{r-1}),$
and thus ${\mathcal{H}}^{r-1}[V({\mathcal{H}})\backslash(U\cup V(F^{\prime}))]$ contains a copy of $S_{2}^{r-1}$. Note that $|V(S_{2}^{r-1})|=2r-3$. For any fixed large constant $2\leq N=o(n)$, by (A.4) and Theorem 1.3, we have
$e({\mathcal{H}}^{r-1})-\sum_{i=2}^{k}\big((r-1)\ell_{i}+1\big)\cdot\binom{n-|U|-1}{r-2}-(N-1)(2r-3)\binom{n-|U|-1}{r-2}$
$\geq$ $|\mathcal{S}_{U}|-\sum_{i=2}^{k}\big((r-1)\ell_{i}+1\big)\cdot\binom{n-|U|-1}{r-2}-o(n)\cdot(2r-3)\binom{n-|U|-1}{r-2}$
$=$ $\Omega(n^{r-1})$
$>$ ${\rm{ex}}_{r-1}(n-|U|,S_{2}^{r-1}).$
This implies that ${\mathcal{H}}^{r-1}[V({\mathcal{H}})\backslash(U\cup V(F^{\prime}))]$ contains a copy of $NS_{2}^{r-1}$. We denote it and its hyperedges by $S^{(1)}\cup\cdots\cup S^{(N)}$ and $e^{(1)}_{1},e^{(1)}_{2},\dots,e^{(N)}_{1},e^{(N)}_{2}$, where $e^{(i)}_{1},e^{(i)}_{2}\in E(S^{(i)})$.
Let $U=\{v_{1},\dots,v_{\lfloor\frac{\ell_{1}+1}{2}\rfloor}\}$. Then we may construct a copy of $P^{r}_{2\lfloor\frac{\ell_{1}+1}{2}\rfloor}$ and a copy of $C^{r}_{2\lfloor\frac{\ell_{1}+1}{2}\rfloor}$ that are vertex-disjoint from $V(F^{\prime})$ as follows.
$P^{r}_{2\lfloor\frac{\ell_{1}+1}{2}\rfloor}:\ \{v_{1}\}\cup e_{2}^{(1)},\{v_{1}\}\cup e_{1}^{(2)},\{v_{2}\}\cup e_{2}^{(2)},\dots,\{v_{\lfloor\frac{\ell_{1}+1}{2}\rfloor}\}\cup e_{2}^{({\lfloor\frac{\ell_{1}+1}{2}\rfloor})},\{v_{\lfloor\frac{\ell_{1}+1}{2}\rfloor}\}\cup e_{1}^{({\lfloor\frac{\ell_{1}+1}{2}\rfloor}+1)},$
$C^{r}_{2\lfloor\frac{\ell_{1}+1}{2}\rfloor}:\ \{v_{1}\}\cup e_{1}^{(1)},\{v_{2}\}\cup e_{2}^{(1)},\{v_{2}\}\cup e_{1}^{(2)},\dots,\{v_{\lfloor\frac{\ell_{1}+1}{2}\rfloor}\}\cup e_{1}^{({\lfloor\frac{\ell_{1}+1}{2}\rfloor})},\{v_{1}\}\cup e_{2}^{({\lfloor\frac{\ell_{1}+1}{2}\rfloor})}.$
Observe that $P^{r}_{\ell_{1}}\subseteq P^{r}_{2\lfloor\frac{\ell_{1}+1}{2}\rfloor}$, and $C^{r}_{2\lfloor\frac{\ell_{1}+1}{2}\rfloor}=C^{r}_{\ell_{1}}$ when $\ell_{1}$ is even. Now we construct a $C^{r}_{\ell_{1}}$ in the case when $\ell_{1}$ is odd. Similarly, by (A.4) and Theorem A.1, we have
$e({\mathcal{H}}^{r-1})-\sum_{i=2}^{k}\big((r-1)\ell_{i}+1\big)\cdot\binom{n-|U|-1}{r-2}-\ell_{1}(2r-3)\binom{n-|U|-1}{r-2}$
$=$ $\Omega(n^{r-1})$
$>$ ${\rm{ex}}_{r-1}(n-|U|,P_{3}^{r-1}).$
This implies that ${\mathcal{H}}^{r-1}[V({\mathcal{H}})\backslash(U\cup V(F^{\prime}))]$ contains a copy of $\ell_{1}S_{2}^{r-1}\cup P_{3}^{r-1}$. We take a copy of $\frac{\ell_{1}-3}{2}S_{2}^{r-1}\cup P_{3}^{r-1}$, and denote it and its hyperedges by $S^{(1)}\cup\cdots\cup S^{(\frac{\ell_{1}-3}{2})}\cup P$ and $e^{(1)}_{1},e^{(1)}_{2},\dots,e^{(\frac{\ell_{1}-3}{2})}_{1},e^{(\frac{\ell_{1}-3}{2})}_{2},e_{1},e_{2},e_{3}$, where $e^{(i)}_{1},e^{(i)}_{2}\in E(S^{(i)})$ and $e_{1},e_{2},e_{3}\in E(P)$. Then we may construct a copy of $C^{r}_{\ell_{1}}$ that are vertex-disjoint from $V(F^{\prime})$ as follows.
$C^{r}_{\ell_{1}}:\ \{v_{1}\}\cup e_{1}^{(1)},\{v_{2}\}\cup e_{2}^{(1)},\dots,\{v_{\frac{\ell_{1}-1}{2}}\}\cup e_{2}^{({\frac{\ell_{1}-3}{2}})},\{v_{\frac{\ell_{1}-1}{2}}\}\cup e_{1},\{v_{\frac{\ell_{1}+1}{2}}\}\cup e_{2},\{v_{1}\}\cup e_{3}.$
Thus, we may always find a copy of $F^{r}_{\ell_{1}}$ that is vertex-disjoint from $V(F^{\prime})$ in $\mathcal{H}$}. This completes the proof. ∎